### v1

# Towards Robust Zero-Shot Reinforcement Learning

Abstract
The recent development of zero-shot reinforcement learning (RL) has opened a new avenue for learning pre-trained generalist policies that can adapt to arbitrary new tasks in a zero-shot manner. While the popular Forward-Backward representations (FB) and related methods have shown promise in zero-shot RL, we empirically found that their modeling lacks expressivity and that extrapolation errors caused by out-of-distribution (OOD) actions during offline learning sometimes lead to biased representations, ultimately resulting in suboptimal performance. To address these issues, we propose Behavior-REgularizEd Zero-shot RL with Expressivity enhancement (BREEZE), an upgraded FB-based framework that simultaneously enhances learning stability, policy extraction capability, and representation learning quality. BREEZE introduces behavioral regularization in zero-shot RL policy learning, transforming policy optimization into a stable in-sample learning paradigm. Additionally, BREEZE extracts the policy using a task-conditioned diffusion model, enabling the generation of high-quality and multimodal action distributions in zero-shot RL settings. Moreover, BREEZE employs expressive attention-based architectures for representation modeling to capture the complex relationships between environmental dynamics. Extensive experiments on ExORL and D4RL Kitchen demonstrate that BREEZE achieves the best or near-the-best performance while exhibiting superior robustness compared to prior offline zero-shot RL methods. The official implementation is available at: https://github.com/Whiterrrrr/BREEZE.
1 Introduction
Reinforcement learning (RL) has become a cornerstone of artificial intelligence, enabling transformative advances in robotics [53], autonomous systems [29], industrial control [71], and large language models (LLM) [60]. However, its real-world adoption faces two persistent challenges: the reliance on human-provided reward functions and its task-specific learning paradigm, which limits adaptability to novel or multiple tasks. These challenges sparked growing interests in zero-shot RL [57, 58, 44, 26, 4, 23, 56, 13], which enables learning a versatile agent through pretraining on reward-free transitions, and then zero-shot adapting to arbitrary reward functions during inference. This opens up new possibilities for developing general-purpose RL agents capable of generalizing across diverse tasks in open-world scenarios.
Existing zero-shot RL approaches mainly fall into two categories: task/skill-conditioned RL [13, 23] and dynamic representation-based methods [4, 3, 58, 26, 56, 44]. The first category encodes demonstrations or reward functions into embeddings as conditioning signals for policy learning. This preserves the task generalization capability of policies, but often results in heavy, manually designed pretraining without optimality guarantees. The second category of methods adopt a more principled approach by decomposing the problem into dynamic representations that can be recomposed for novel tasks without retraining. Among dynamic representation-based methods, Forward-Backward representations (FB) [57, 58] have recently attracted notable attention, factorizing occupancy measures into two components: a forward representation that captures policy dynamics and a backward representation that encodes global state information. Through offline, unsupervised pretraining, the FB framework learns linearized representations that approximate value functions for arbitrary tasks, holding great promise for zero-shot generalization.
However, despite the elegant theoretical framework provided by the FB representations, our empirical studies reveal that the successor measures learned through existing FB-based methods are often inconsistent and biased, thereby compromising the stability and overall performance (see Section 3 for details). In this paper, we identify two causes for the shortcomings of existing FB-based methods. First, learning complex approximators and multimodal behaviors requires highly expressive models, which current FB methods lack in both their representations and policy. Second, the offline, unsupervised pretraining stage suffers from extrapolation error due to out-of-distribution (OOD) actions, which is similar to offline RL [17, 32, 33], but exhibits more complex behavior. As our results show, naively integrating value constraints can be ineffective [26], indicating a need for a more delicate OOD regularization mechanism.
Based on the above observations, we propose Behavior-REgularizEd Zero-shot RL with Expressivity enhancement (BREEZE), a novel FB-based framework that simultaneously enhances offline learning stability and zero-shot generalization capability. First, we introduce a behavior-regularized reformulation of FB, which mitigates extrapolation errors while preserving the fidelity of representations. Second, we extract the policy using a task-conditioned diffusion model, enabling high-quality multimodal action distributions in zero-shot RL settings. Finally, we employ expressive representation networks based on the attention architecture to capture complicated dynamics. We conducted extensive experiments on the ExORL [70] benchmark and the D4RL [14] Kitchen dataset, under both full datasets and small-sample data regimes. The results demonstrate that BREEZE achieves the best or near-the-best performance while exhibiting superior robustness.
2 Preliminary
Reward-free Markov decision process (MDP).
Zero-shot RL is typically formulated as a reward-free Markov Decision Process (MDP) [51], defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\gamma)$, consisting of a state space $\mathcal{S}$, an action space $\mathcal{A}$, a transition kernel $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, and a discount factor $\gamma\in(0,1)$. Given an initial state-action pair $(s_{0},a_{0})\in\mathcal{S}\times\mathcal{A}$ and a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$, we denote $\text{Pr}(\cdot|s_{0},a_{0},\pi)$ a probability and $\mathbb{E}[\cdot|s_{0},a_{0},\pi]$ an expectation under the state-action trajectories $(s_{0},a_{0},...,s_{t},a_{t})_{t\geq 0}$ generated by sampling $s_{t}\sim\mathcal{P}(\cdot|s_{t-1},a_{t-1})$ and $a_{t}\sim\pi(\cdot|s_{t})$. The state transition under $\pi$ is given by $\mathcal{P}_{\pi}(\mathrm{d}s^{\prime}|s)=\int\mathcal{P}(\mathrm{d}s^{\prime}|s,a)\pi(\mathrm{d}a|s)$.
Approximate dynamic programming-based RL [47] uses an action-value $Q$-function, or optionally a state-value $V$-function. The $Q$-function and $V$-function for $\pi$ starting at $s_{0},a_{0}$ under a given reward function $r:\mathcal{S}\to\mathbb{R}$ are respectively defined as $Q^{\pi}_{r}(s_{0},a_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},a_{0},\pi]$ and $V^{\pi}_{r}(s_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},\pi]$. The goal of the zero-shot RL problem is to train a task-agnostic policy given offline dataset $\mathcal{D}=\{s_{i},a_{i},s_{i+1}\}_{i=1}^{N}$ generated by an unknown behavior policy $\mu(\cdot|s)$, that can later generalize to any downstream task $z$ defined by a reward function $r_{\text{eval}}:\mathcal{S}\to\mathbb{R}$, i.e. find $\max_{\pi}\ \mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{\text{eval}}\left(s_{t+1}\right)|s_{0},a_{0},\pi]$, relying only on a small set of reward-labeled samples with no further finetuning.
Forward-Backward representations (FB).
FB introduces a rank-$d$ approximation of the successor measure, which is defined as the expected discounted occurrences of future states $s_{+}\in\mathcal{S}_{+}$ after starting from $(s_{0},a_{0})$ under a policy $\pi$:
$M^{\pi}(s_{0},a_{0},\mathcal{S}_{+}):=\sum_{t\geq 0}\gamma^{t}\,\text{Pr}\big(s_{t+1}\in\mathcal{S}_{+}\mid s_{0},a_{0},\pi\big)\,\quad\forall\mathcal{S}_{+}\subset\mathcal{S}.$ (1)
Following this definition, $Q$-function under policy $\pi$ can be expressed as:
$Q_{r}^{\pi}(s,a)=\int_{s_{+}\in\mathcal{S}}M^{\pi}(s,a,ds_{+})\,r(s_{+}).$ (2)
Given representation space $\mathbb{R}^{d}$, state distribution $\rho$, task vector $z\in\mathbb{R}^{d}$ and policy $\pi_{z}$ parameterized by $z$, FB decouples $M^{\pi_{z}}$ by a pair of forward representation $F:\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and backward representation $B:\mathcal{S}\to\mathbb{R}^{d}$, resulting optimal policy $\pi_{z}$ as:
$M^{\pi_{z}}(s_{0},a_{0},ds_{+})\approx F(s_{0},a_{0},z)^{\top}B(s_{+})\rho(ds_{+}),\quad\pi_{z}(s):=\arg\max_{a}F(s,a,z)^{\top}z,$ (3)
where $z$ is defined as $z:=\mathbb{E}_{s\sim\rho}[r(s)B(s)]$ from a few samples with a known reward function $r$. The corresponding $Q$-function can be obtained immediately as $Q_{z}(s,a)=F(s,a,z)^{\top}z$.
Since the successor measure $M^{\pi_{z}}$ satisfies a Bellman-like equation $M^{\pi_{z}}=\mathcal{P}+\gamma\mathcal{P}_{\pi_{z}}M^{\pi_{z}}$, FB derives the temporal difference (TD) objective [50] on $M^{\pi_{z}}$ as:
$\begin{gathered}\mathcal{L}_{\text{FB}}=\mathbb{E}_{\begin{subarray}{c}(s_{t},a_{t},s_{t+1})\sim D\\ s_{+}\sim D\end{subarray}}\left[\left(F(s_{t},a_{t},z)^{\top}B(s_{+})-\gamma\bar{F}(s_{t+1},\pi_{z}(s_{t+1}),z)^{\top}\bar{B}(s_{+})\right)^{2}\right]\\ -2\,\mathbb{E}_{(s_{t},a_{t},s_{t+1})\sim D}\left[F(s_{t},a_{t},z)^{\top}B(s_{t+1})\right],\end{gathered}$ (4)
and incorporate the following objective on $F$ to enforce the Bellman property on the $Q$-function:
[…]
where $B(s_{t+1})^{\top}\mathbb{E}_{D}[BB^{\top}]^{-1}z$ is the implicit reward estimation and $D$ denote the data distribution.
Recent extensions of FB incorporate additional regularization to improve learning performance. For example, Jeen et al. [26] noted that the use of $\pi_{z}(s_{t+1})$ in Eq. (4) during offline learning could introduce OOD extrapolation errors. To resolve this problem, they introduce additional CQL-style [33] regularizers on $M^{\pi_{z}}$ (MCFB) or $Q_{z}$ (VCFB) in $\mathcal{L}_{\text{FB}}$ for OOD regularization (i.e., push down the corresponding values on OOD actions and push up values evaluated on offline data samples).
3 Pitfalls of Exiting FB-Based Methods
(a) Distribution of empirically evaluated $M^{\pi_{z}}$
(b) Distribution of empirically evaluated $Q_{z}$
Figure 1: Visualization of the empirical $M^{\pi_{z}}$ and $Q_{z}$ distributions during evaluation stage. We conduct experiments on FB-based methods on two Walker tasks using ExORL [70] RND [5] dataset. We use the learned $F$, $B$ representations to recover the task vector $z$ and compute $M^{\pi_{z}}=F^{\top}B$ and $Q_{z}=F^{\top}z$. Both the vanilla FB [58] and MCFB [26] result in a proportion of error scaling values.
FB offers an elegant theoretical foundation for zero-shot RL; however, our empirical findings reveal that FB-based methods often produce inconsistent and biased representations in practice. To illustrate this, we evaluate the $M^{\pi_{z}}$- and $Q_{z}$-value distributions derived from the learned $F$ and $B$ representations of the vanilla FB [58], the more recent MCFB [26], and our method on ExORL [70] benchmark with RND [5] datasets. For each method, we use its $B$ representation to derive the task vector $z$ under default settings with a fixed number of transitions. The $M^{\pi_{z}}$ values are computed from episode-start state-action pairs $(s_{0},a_{0})\sim\rho_{0}$ and randomly sampled transitions $s_{+}\sim\mathcal{D}$, while $Q$-values are evaluated using policy-induced actions $a\sim\pi_{z}(s)$. Results across additional environments and with method VCFB [26] are provided in Appendix D.3.
Figure 1 illustrates the value distributions from two tasks in the Walker domain. Although the successor measure $M^{\pi_{z}}$ is mathematically a positive quantity representing future state occupancy, the representations learned by existing FB-based methods fail to accurately capture this property. These distributions exhibit two notable discrepancies: a scale mismatch characterized by enormous absolute values, and the presence of a considerable number of invalid negative values. Such inaccuracies in the $F$ and $B$ representations subsequently affect downstream $Q$-value estimates, creating a consistent prediction bias across the system, which is further worsened by the suboptimal action rollout. The CQL-style regularizer in MCFB partially mitigates the issue, shifting the distributions of $M^{\pi_{z}}$ and $Q_{z}$ toward more plausible ranges. However, this correction is incomplete, and significant estimation bias persists, indicating the need for a more effective regularization strategy. In comparison, our method generates distributions for $M^{\pi_{z}}$- and $Q_{z}$-value that more closely match theoretical expectations, and can also result in a high-quality policy.
(a) Impact of $F$ Networks
(b) Impact of $B$ Networks
Figure 2: Investigations of the modeling architecture on the vanilla FB [58]. We utilize the peak observed in the training stage of each run as a proxy for the architecture’s capability ceiling.
The challenges in estimating $M^{\pi_{z}}$- and $Q_{z}$-value can be primarily attributed to the limited supervisory signal in the vanilla FB learning objective (Eq. (4)), which fails to constrain the scale of the learned representations adequately. Additionally, modeling successor measures and zero-shot policies across all possible task vectors $z$ introduces substantial complexity, necessitating models with high expressive capacity. While the marginal gains from MCFB’s regularizer suggest its potential utility, a more effective formulation appears necessary. These observations are further corroborated by the results from our method, which integrates a refined OOD regularization strategy and employs a more expressive model architecture. As a result, the distributions of $M^{\pi_{z}}$ and $Q_{z}$ values are primarily confined to a reasonable range. Importantly, as illustrated in Figure 2, these improvements are not attributable to increased model capacity alone: simply scaling up the original MLP-based FB networks does not yield measurable gains. In contrast, the architectural modifications designed in this work lead to a marked improvement in performance, underscoring the influence of model expressivity on the effectiveness of zero-shot learning.
4 Behavior-Regularized Zero-Shot RL with Expressivity Enhancement
Motivated by these findings, we propose BREEZE, a framework designed to mitigate the OOD and expressivity issues in FB-based methods. We begin by imposing a behavioral regularizer on the offline zero-shot RL problem, followed by the transformation of the learning process into an in-sample weighted regression paradigm, which can naturally integrate powerful generative models to enhance policy capability. Recognizing the importance of expressivity for learning arbitrary tasks, we describe our practical architecture for stronger representation modeling.
4.1 Behavior-regularized Optimization
The main OOD extrapolation issue in the vanilla FB learning losses $\mathcal{L}_{FB}$ and $\mathcal{L}_{F}$ in Eq. (4) and Eq. (5) lies in the policy-generated actions $\hat{a}_{t+1}\sim\pi_{z}(s_{t+1})$ in $F(s_{t+1},\pi_{z}(s_{t+1}),z)$ during offline learning. Since $\pi_{z}$ is trained to maximize $F(s,a,z)^{\top}z$, it may produce actions that appear optimal under the learned representation but are actually OOD and overestimated, leading to extrapolation errors. Our solution to this is to introduce two designs: 1) stabilize the $F(s_{t+1},\pi_{z}(s_{t+1}),z)$ value estimate in the $\mathcal{L}_{FB}$; 2) ensure that $\pi_{z}$ is better regularized by the dataset samples during policy optimization.
Behavior-regularized representation guidance.
To stabilize the learning of representation learning, we consider a regularized version of $\mathcal{L}_{F}$ in Eq. (5) by introducing the task-conditioned state-value function $V$ with respect to policy $\pi$, given by:
$\begin{gathered}V_{\pi_{z}}(s,z):={\max_{\begin{subarray}{c}a\in A\\ \mathrm{s.t.}\,\mu(a|s)>0\end{subarray}}}F(s,a,z)^{\top}z,\end{gathered}$ (6)
where $z$ is the task vector and $\mu(a|s)$ is the behavior policy in dataset $\mathcal{D}$. Compared with the traditional formulation of state-value function, this optimization problem aims to learn an optimal task-conditioned $V$-function solely from dataset samples without explicitly estimating $\mu(a|s)$. It can be solved through a class of value objectives according to different regularization formulations [67, 18, 30]. In this work, we leverage the commonly used expectile regression as in IQL [30]:
$\mathcal{L}_{V_{\pi_{z}}}=\mathbb{E}_{(s,a)\sim\mathcal{D},z\sim\mathcal{Z}}\left[L_{2}^{\tau}\left(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)\right)\right],$ (7)
where $L_{2}^{\tau}(u)=|\tau-\mathbb{I}(u>0)|u^{2}$, with $\tau>0.5$ serving as the expectile parameter. By minimizing the above equations, we can obtain a well-regularized and nearly optimal state-value function $V_{\pi_{z}}$ for different tasks. Given the reward function $r_{z}$, associated with task $z$, we have the following modified policy evaluation operator $\mathcal{T}^{\pi}$ given by
$\begin{gathered}(\mathcal{T}^{\pi}Q_{z})(s,a,z):=\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s^{\prime}|s,a)}\left[r_{z}(s^{\prime})+\gamma V_{\pi_{z}}(s^{\prime},z)\right],\end{gathered}$ (8)
This formulation of $\mathcal{T}^{\pi}$ implicitly introduces regularization to the $Q$-function. Following above, we modified $\mathcal{L}_{F}$ in Eq. (5) as following $\mathcal{L}_{F\text{-reg}}$:
$\mathcal{L}_{F\text{-reg}}=\mathbb{E}_{\begin{subarray}{c}(s,a,s^{\prime})\sim\mathcal{D}\end{subarray}}\left[\left(F(s,a,z)^{\top}z-B(s^{\prime})^{\top}\mathbb{E}_{D}[BB^{\top}]^{-1}z-\gamma V_{\pi_{z}}(s_{t+1},z)\right)^{2}\right].$
By substituting the potentially unstable target $Q$-approximation $F(s_{t+1},\pi_{z}(s_{t+1}),z)^{\top}z$ with a more well-behaved state-value function $V_{\pi_{z}}$, we can greatly stabilize the learning of representation, as well as respond to optimality demands simultaneously. In comparison with the direct constraint on unseen value approximation, the behavioral regularization guidance ensures maximum preservation of the representation structure, while being flexible for tuning the degree of conservatism.
Behavior-regularized policy extraction.
Having introduced regularization for the value function, we now turn to policy learning, which prior work suggests can be equally or even more critical for final performance [43]. A good policy should effectively leverage the high-quality behavior and further generalize near data distribution. To ensure that the policy is well regularized by the dataset samples, we replace the policy objective in Eq. (3) with the following behavior-regularized optimization problem:
$\max_{\pi_{z}}\mathbb{E}_{a\sim\pi_{z}(\cdot|s)}\left[F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)\right]\quad\text{s.t.}\ \int_{a}\pi_{z}(s)da=1,\,\forall s\in\mathcal{S},\quad D_{\text{KL}}(\pi_{z}\|\mu)\leq\epsilon,$ (10)
where the term $F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)$ acts as an advantage function maximization, which is equivalent to maximizing the $Q$-function. The KL constraint anchors $\pi_{z}$ to the behavior policy $\mu$, thereby preventing distributional shift in the offline setting. Given the optimal representation $F$ and $V_{\pi_{z}}$, we can obtain the closed-form solution for the constrained optimization objective Eq. (10) by deriving the Lagrangian objective with respect to the policy and setting its derivative to zero, as shown in Proposition 1 below:
Proposition 1.
The solution to the constrained optimization problem in Eq. (10) yields an optimal policy of the form (See Appendix A for proof):
$\pi_{z}^{*}(s)\propto\mu(a|s)\exp\left(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\right),$ (11)
where $1/\alpha$ is the Lagrangian multiplier for the KL constraint.
$\alpha$ acts as a temperature that controls the balance between the regularization strength of the behavioral policy and the optimization of the value functions. This closed-form solution elegantly combines the distribution of the behavior policy with the optimal value functions, encouraging the policy to favor high-advantage actions while ensuring all actions remain within the support of the dataset, thus balancing performance with OOD avoidance. With this stable guarantee, we now aim to enhance the policy generalization capability in the next section.
4.2 Policy Extraction via Task-Conditioned Diffusion Model
To compute the solution of the weighted behavior cloning objective in Eq. (11) in practice, a key challenge arises: how to efficiently model and sample from potentially highly complex and diverse distributions distilled from a reweighted behavior policy?
Direct weighted behavior cloning with a Gaussian policy often fails to capture the complex, multi-modal policy distributions [62, 20, 7], which is a demand for arbitrary task learning. This motivates the use of diffusion models [48, 22, 37], known for their capacity to learn complex distributions through iterative denoising. Rather than relying on guidance-based techniques that introduce a separate, time-dependent term to steer the sampling process, we draw inspiration from recent advances in weighted regression for diffusion models [7, 20, 27, 72]. Building on the theoretical foundation of prior work [72], we formalize in Proposition 2 how the optimal policy $\pi_{z}^{*}$ can be extracted using a diffusion model trained with a weighted regression objective:
Proposition 2 (Task-conditioned diffusion policy extraction via weighted regression).
The extraction of optimal policy $\pi_{z}^{*}$ in Eq. (11) can be achieved by (i) minimizing the weighted regression loss defined as:
$\min_{\theta}\mathbb{E}_{t\sim\mathcal{U}([0,T]),\epsilon\sim\mathcal{N}(0,I),(s,a)\sim\mathcal{D}}\Big[\exp\Big(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\Big)\|\epsilon-\epsilon_{\theta,z}(a_{t},s,z,t)\|_{2}^{2}\Big],$ (12)
with expectations taken over $t\sim\mathcal{U}([0,T])$, $\epsilon\sim\mathcal{N}(0,I)$, and $(s,a)\sim\mathcal{D}$, where $a_{t}=\alpha_{t}a+\sigma_{t}\epsilon$ follows the forward process $\mathcal{N}(a_{t}|\alpha_{t}a,\sigma_{t}^{2}I)$ parameterized by noise schedules $\alpha_{t},\sigma_{t}$ and $z\sim\mathbb{R}^{d}$ denotes the corresponding task; and (ii) sampling from $\pi_{z}^{*}$ by solving the corresponding diffusion ODEs/SDEs with the learned $\epsilon_{\theta,z}$. (See Appendix A for the detailed proof.)
With the above objective, we can avoid the need for learning the additional time-dependent guidance term and thereby reduce the complexity, while getting a stable policy.
For action selection, we employ a rejection sampling mechanism to boost policy performance. Specifically, we first sample $K$ candidate actions $\{a^{(1)},\ldots,a^{(K)}\}\sim\pi_{z}(s)$ through the policy rollout. We then evaluate each candidate using the $Q$-function (approximated by $F(s,a,z)^{\top}z$) and select the action with the highest value:
$a^{\star}\triangleq\arg\max_{a\in\{a^{(1)},\ldots,a^{(K)}\sim\pi_{z}(s)\}}F(s,a,z)^{\top}z.$ (13)
This two-stage approach ensures the policy balances both conservatism and optimism. The diffusion model generates diverse, in-distribution candidates, while the selection step identifies the action with the highest expected return. This combination of expressive generative modeling and value-based selection is crucial for achieving robust, high-performance zero-shot generalization.
4.3 Expressivity Enhancement for Representations Modeling
The effectiveness of any policy—particularly a highly expressive diffusion policy—in leveraging learned action weights hinges on accurate value estimation. Biased value estimates can degrade policy learning to simple imitation, failing to redistribute probability mass toward superior actions. To fully exploit the representational capacity of our diffusion policy, the underlying value representations must precisely capture complex task and dynamics relationships. We thus introduce enhanced network architectures for the forward ($F$) and backward ($B$) representations.
Figure 3: The architectures in BREEZE
As shown in Section 3, we empirically identify a pair of attention-based architectures that improve zero-shot performance over the original FB implementations. Below, we detail the design of our networks, illustrated in Figure 3.
Forward Network.
The forward network encodes state-task and state-action pairs using two separate linear encoders, following Touati et al. [58]. The state-task encoder captures task-conditioned dynamics, while the state-action encoder extracts agent behavior patterns. These two feature sets correspond to semantically distinct concepts in the MDP. To accurately approximate the future occurrence measure $M(s,a,s^{\prime},z)$ in Eq. 3, the representation $F$ must integrate information from both behavioral and task contexts, capturing their interdependencies effectively. We therefore model the two encoded features as a length-2 sequence and process them through self-attention blocks. This allows bidirectional feature refinement between task conditions and agent behaviors. The resulting representation is projected to a $d$-dimensional space via a linear layer.
Backward Network.
As defined in Eq. 3, the backward representation $B$ encodes state-level structural information, serving as a global embedding of the environment. Intuitively, $B$ should employ a more complex architecture, maintaining orthogonality and enforcing $F$ alignment. We equip $B$ with a stack of standard transformer-based networks with multi-head attention. The final output is projected to a $d$-dimensional space, consistent with the forward network.
These architecturally enhanced representations enable BREEZE to model complex relationships more accurately, leading to improved value estimates and policy performance. Further implementation details are provided in Appendix B.4.
Table 1: IQM results on ExORL[70] benchamrk. We report the best overall performance when all tasks perform well. Each value is averaged over 10 rollouts on 5 random seeds.
Dataset Domain SF-LAP FB VCFB MCFB HILP BREEZE
RND Walker $516\pm 97$ $661\pm 10$ $653\pm 22$ $59\pm 51 66$$65\pm 33$ $693\pm 16$
Jaco $18\pm 18$ $32\pm 23$ $46\pm 35$ $41\pm 34$ $52\pm 21$ $84\pm 14$
Quadruped $330\pm 165$ $671\pm 14$ $609\pm 29$ $684\pm 18$ $674\pm 28$ $725\pm 23$
APS Walker $324\pm 24$ $517\pm 99$ $487\pm 75$ $578\pm 35$ $643\pm 22$ $637\pm 21$
Jaco $39\pm 26$ $22\pm 14$ $0\pm 18 22$$2\pm 3$ $84\pm 16$ $132\pm 16$
Quadruped $498\pm 160$ $68\pm 29 66$$64\pm 3$ $659\pm 50$ $679\pm 14$ $698\pm 24$
PROTO Walker $382\pm 129$ $650\pm 19$ $611\pm 94$ $602\pm 112$ $715\pm 31$ $663\pm 19$
Jaco $15\pm 14$ $21\pm 26$ $13\pm 12$ $20\pm 21$ $44\pm 19$ $74\pm 26$
Quadruped $199\pm 10$ $222\pm 107$ $185\pm 72$ $219\pm 135$ $216\pm 54$ $389\pm 44$
DIAYN Walker $239\pm 79$ $338\pm 74$ $268\pm 67$ $268\pm 97$ $461\pm 64$ $463\pm 42$
Jaco $32\pm 26$ $22\pm 6$ $24\pm 3$ $15\pm 1$ $52\pm 7$ $78\pm 11$
Quadruped $207\pm 168$ $562\pm 23$ $511\pm 37$ $643\pm 14$ $70\pm 4 66$$66\pm 2$
Table 2: IQM results with 100k-dataset. Experiments on randomly sampled 100k-transition data from each dataset on the ExORL benchmark
Dataset Domain FB VCFB MCFB BREEZE (ours)
RND Walker $264\pm 33$ $350\pm 29$ $287\pm 48$ $525\pm 13$
Jaco $7\pm 5$ $9\pm 2$ $13\pm 7$ $36\pm 5$
Quadruped $176\pm 123$ $233\pm 52$ $123\pm 61$ $474\pm 21$
APS Walker $370\pm 66$ $416\pm 10$ $389\pm 77$ $539\pm 15$
Jaco $21\pm 17$ $14\pm 13$ $29\pm 27$ $38\pm 9$
Quadruped $340\pm 29$ $351\pm 57$ $318\pm 122$ $556\pm 52$
PROTO Walker $415\pm 19$ $513\pm 31$ $463\pm 11$ $553\pm 18$
Jaco $16\pm 2$ $18\pm 12$ $12\pm 7$ $29\pm 12$
Quadruped $198\pm 111$ $106\pm 103$ $240\pm 134$ $181\pm 60$
DIAYN Walker $202\pm 94$ $210\pm 81$ $196\pm 26$ $330\pm 43$
Jaco $17\pm 11$ $18\pm 5$ $0\pm 26 22$$2\pm 15$
Quadruped $295\pm 46$ $288\pm 48$ $286\pm 34$ $446\pm 78$
Table 3: Normalized scores on Kitchen environment
5 Experiments
We conduct extensive evaluations of BREEZE against previous offline zero-shot RL algorithms across various challenging domains with distinct tasks. We also present the ablations on component choices and hyperparameters.
Environmental setup.
Our main experiments are conducted on the ExORL [70] benchmark, which provides a variety of datasets collected by several unsupervised RL algorithms [34]. We select datasets collected by 4 algorithms: RND [5], APS [38], DIAYN [11], PROTO [69]. The experiments span 3 domains and 12 tasks (Walker: Stand, Walk, Run, Flip; Jaco: Reach Top/Bottom Left/Right; Quadruped: Stand, Walk, Run, Jump), bringing the total to 48 state-based complex tasks for performance calculation. In addition, we consider four challenging multi-stage tasks in the D4RL [14] Franka Kitchen [19] domain with two datasets (mixed and partial), which require long-horizon sequential manipulation on 4 subtasks. The overall setups on two benchmarks involve 2 locomotion domains and 2 manipulation (goal-reaching) domains. For all goal-conditioned domains, we use the backward representation to calculate the goal features as the task vector $z$.
Baselines.
We only consider offline algorithms that could perform a zero-shot policy generalization for our evaluation. Our baselines includes: 1) SF-LAP [3, 4]: successor feature-based method with basic features to be Laplacian Eigenfunction (LAP) [64], 2) vanilla FB [57], 3) VCFB and MCFB [26]: FB-based offline algorithms with CQL [33]-style regularizer on successor measure $M^{\pi}$ and $Q$-value separately for OOD issue avoidance, 4) HILP [44]: a state-based representation modeling algorithm keeping distance-preserving temporal structures in latent space having zero-shot capability.
Evaluation.
Our experiments are designed to evaluate two key aspects: zero-shot performance and robustness. To assess the overall zero-shot capability, we experiment on the full transitions of each dataset on ExORL and D4RL Kitchen. For robustness, we first present performance curves across training time to demonstrate learning stability and then conduct experiments on a uniformly subsampled 100,000 transitions to examine performance in a limited and low-quality data coverage scenario, following the evaluation process of Jeen et al. [26]. We record the Interquartile Mean (IQM), a robust metric that reduces the influence of outliers, in the training stage. We evaluate the checkpoints every 10,000 updates with 10 rollouts for each experiment. On Kitchen, we multiply the returns by 25 for normalization, following Park et al. [44]. Further implementation and experimental details are provided in Appendix B.
5.1 Experimental Results
We analyze the findings of our experiments through the following key questions:
  • Does BREEZE possess zero-shot capability to novel tasks? Following previous studies [1, 26], we record the highest aggregated mean performance across random seeds in the offline pretrain stage. Table 3 reports this score across all tasks on each domain of ExORL, averaged on 5 random seeds, with standard deviations. BREEZE achieves the best or near-best returns in most domains, demonstrating superior zero-shot generalization capability. Moreover, BREEZE can enhance vanilla FB’s generalization capability in long-horizon tasks. As shown in Table 3, we draw the box with the performance reported in the previous study [44], and record BREEZE’s top averaged returns across 4 random seeds. While vanilla FB struggles in these long-horizon tasks, BREEZE achieves significantly higher performance, suggesting its ability to better leverage the dataset for complex sequential manipulation.
    More results are provided in Appendix D.2.
  • How does BREEZE quantitatively compare to baselines in terms of stability and convergence speed? We provide the learning curves in RND [5] datasets as Figure 4. In locomotion domains (i.e., Quadruped and Walker), BREEZE converges faster to higher performance, with smoother curves and lower variance (as shown by the small shaded area). In the manipulation domain (i.e., Jaco), BREEZE substantially outperforms all baselines with a higher learning speed. Overall, BREEZE demonstrates faster convergence and enhanced stability across both locomotion and manipulation scenarios.
  • How does BREEZE perform facing different quality, diversity of datasets? Table 3 shows results on 100,000-transition subsets of ExORL datasets. BREEZE maintains a pronounced advantage over FB-based baselines in this small-sample regime. Compared to explicit constraint methods (MCFB/VCFB), these results highlight the importance of behavior alignment across different data regimes. BREEZE also exhibits superior stability, as shown in Figure 5.
Refer to Appendix D.1 for more learning curves.
Figure 4: Learning Curves on ExORL RND. The solid lines represent the average return over 5 random seeds, while the shaded area denotes the standard deviation.
Figure 5: Learning Curves on 100k-subsample transitions on ExORL APS Walker. The solid lines represent the average return over 3 random seeds, while the shaded area denotes the standard deviation.
Dataset w/o FB Enhancement w/o Diffusion BREEZE
[…]
Table 4: Ablation on necessity of FB enhancement and diffusion policy (3 seeds).
Figure 6: Ablation on the temperature.
Figure 7: Ablation on the expectile factor.
5.2 Ablation Studies
Hyperparameter choices.
We ablate two key hyperparameters in our behavior-constrained algorithm: the expectile value $\tau$ and the temperature $\alpha$. As shown in Figure 6 and Figure 7 (Quadruped domain, RND dataset, 3 random seeds), performance improves near monotonically with $\tau$ and peaks at $\alpha=0.05$, and similarly to other domains. This indicates that, while in-sample conservatism is crucial, the algorithm benefits from more aggressive, value-weighted optimization. Consequently, we select $\tau=0.99$ and $\alpha=0.05$ as defaults in ExORL, optimally balancing conservatism with performance for downstream tasks.
Necessity of each component.
We ablate our two core innovations—an enhanced representation model and a diffusion policy—to assess their necessity in RND and APS datasets. BREEZE is ablated using: a gaussian policy variant (w/o Diffusion) and the baseline architecture variant (w/o FB Enhancement). We process transitions with data normalization for the diffusion policy, as in a commonly used preprocessing step in previous diffusion-based methods, and cancel this process for backing to the default policy setting to have a reasonable comparison on the components’ contributions. As shown in Table 4, each component somehow individually slightly boosts performance, but their synergy creates a leap in zero-shot capability. This demonstrates a critical interdependence: the design of BREEZE is a logical combination to handle the system issue.
6 Related Works
Unsupervised zero-shot RL.
Recent advancements in unsupervised zero-shot RL [34, 44, 13] mainly build upon the utilization of successor representation (SR) [10], with two main branches: successor feature (SF) [3, 4] and forward-backward representation (FB) [57, 58]. Both approaches require decomposing and linearizing reward-aware dynamics to enable zero-shot inference for arbitrary downstream rewards. Other approaches relax the linearity assumption by modeling SR as full distributions using generative models [25, 55, 63, 12]. Among these, FB has emerged as a key framework, reducing the reward dependency by learning the successor measures $M$ instead of $Q$ or $V$, and improving tractability by factorizing $M$ into a forward and backward components. Extensions like MCFB/VCFB [26] build on offline RL adding value constraints to reduce the overestimation from OOD rollouts. FB-CPR [56] introduces regularizes the $Q$-estimation by imitating pre-collected datasets in an online setting. FB-AWARE [6] introduces an autoregressive mechanism to handle the over-linearization problem in vanilla FB. Unlike SR-based approaches, HILP [44] structures the latent space so that distances correspond to transition times between states, enabling hierarchical, goal-conditioned, and zero-shot policy learning. Meanwhile, functional-encoder methods [13, 23] directly encode task information by pretraining on manually designed reward functions, but are constrained by the need for extensive reward engineering and fine-tuning.
Offline RL.
Numerous studies address offline RL, in which learning is restricted to a fixed dataset without online interaction, leading to distributional shift. The earliest approaches use policy constraints, enforcing the learned policy to stay close to behavior policy [17, 32, 15, 35] or the distribution coverage to stay near the dataset [36, 8]. Another established line of work employs value regularization, which directly penalizes the value estimates of OOD actions [33, 31, 42, 68, 66]. Instead of an explicit constraint, in-sample learning methods [30, 67, 65, 61] improve stability by learning values and policies only from state-action pairs within the dataset.
Moving beyond conventional Gaussian policies, recent studies have highlighted the importance of modeling multimodal action distributions [62, 20, 7]. Diffusion models [48, 22, 37], known for their strong multi-modal modeling capability, have been well studied in imitation and trajectory modeling [24, 73, 9, 41, 52], and have recently been integrated into offline RL frameworks. Some approaches use the value function as guidance for sampling [39, 40]. Other works, such as SfBC [7], IDQL [20], and EDP [27], extract a diffusion policy via weighted regression under the in-sample learning framework.
7 Conclusion
In this work, we present BREEZE, a novel framework that mitigates the improper scaling issue in existing FB-based zero-shot RL methods. Specifically, BREEZE tackles two critical limitations: offline extrapolation errors and constrained policy expressivity. Our solution integrates behavior-regularized value estimation with a task-conditioned diffusion policy, enabling stable in-sample learning while capturing complex, multimodal action distributions. Coupled with expressive attention-based representations, BREEZE more accurately models dynamics and value functions. Extensive experiments demonstrate that BREEZE consistently outperforms existing methods across various benchmarks, underscoring the importance of calibrated regularization and sufficient model capacity for zero-shot generalization. While BREEZE demonstrates improved robustness, its primary limitation is the increased computational cost from diffusion-based sampling—a common trade-off for high-performance generative policies. Further discussion is provided in Appendix E.
Acknowledgment
This work is supported by the Wuxi Research Institute of Applied Technologies, Tsinghua University, under Grant 20242001120, and funded by Horizon Robotics, AsiaInfo, and the Xiongan AI Institute.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, the paper discusses its limitations in Appendix E, including the high computational demand.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Yes, the paper provides the full set of assumptions and complete proofs for all theorems in the appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, the paper fully discloses all information needed to reproduce the main experimental results. Pseudocode, all experimental settings, hyperparameters, and baseline details are included in the appendix B-C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Yes, the code will be made available after the open-source approval process is completed.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, all hyperparameters and experimental setup details necessary to understand the results are provided in the appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Yes, for all algorithms, several random seeds are used. All results are presented as mean ± standard deviation.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, the Appendix C and E provide detailed information about the specific computational devices used for the experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We make sure to preserve anonymity.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes, the main paper discusses both potential positive societal impacts and limitations of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, all sources of data and code are properly credited in the appendix B, with licenses and terms of use clearly indicated.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification:
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer:[N/A]
  69. Justification:
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification:
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
Appendix A Theoretical Proofs
A.1 Proof of Proposition 1
Proposition 1.
The closed-form optimal policy for Eq. (10) follows the form:
$\pi_{z}^{*}(s)\propto\mu(a|s)\exp\left(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\right),$ (14)
where $\mu$ denotes the behavior distribution of the dataset.
Proof.
Construct the Lagrangian with dual variables $\lambda_{1}$ and $\lambda_{2}$ for the equality and KL-divergence constraints, respectively:
$\mathcal{L}(\pi_{z},\lambda_{1},\lambda_{2})=\mathbb{E}_{a\sim\pi_{z}}\left[F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)\right]-\lambda_{1}\left(\int\pi_{z}(a|s)\,da-1\right)-\lambda_{2}\left(D_{\text{KL}}(\pi_{z}\|\mu)-\epsilon\right).$ (15)
Taking the functional derivative of $\mathcal{L}$ with respect to $\pi_{z}(a|s)$ and setting it to zero yields:
$\frac{\partial\mathcal{L}}{\partial\pi_{z}}=F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)-\lambda_{2}\left(\log\pi_{z}(a|s)-\log\mu(a|s)+1\right)-\lambda_{1}=0.$ (16)
Solving above for $\pi_{z}$ and absorbing constant terms into the normalization factor $Z(s,z)$ gives:
$\pi_{z}^{*}(a|s)=\frac{\mu(a|s)}{Z(s,z)}\exp\left(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\right).$ (17)
where $\alpha=1/\lambda_{2}$ and the partition function is defined as:
$Z(s,z)=\int\mu(a|s)\exp\left(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\right)\,da,$ (18)
This ensures the policy is properly normalized. The proof is completed. ∎
A.2 Proof of Proposition 2
We begin by conducting a theoretical review of diffusion models, and then demonstrate how to derive the optimal policy through task-conditioned weighted regression.
Diffusion (probabilistic) models. Diffusion models [48, 22, 37] are powerful generative models designed to learn and replicate complex distributions. Given an $N$-sample dataset $\mathcal{D}=\{x_{0}^{i}\}^{N}_{i=1}$ from an unknown distribution $q(x_{0})$, diffusion models can accurately approximate $q(x_{0})$ and generate new samples from the approximation by the following two processes:
  • Forward process. The forward process gradually adds Gaussian noise on sample $x_{0}$ from time 0 to $x_{T}$ over $T$ timesteps, with transformation distribution follows:
    $q_{t0}(x_{t}|x_{0})=\mathcal{N}(x_{t}|\alpha_{t}x_{0},\sigma_{t}^{2}I),$ (19)
    where $\alpha_{t}$ and $\sigma_{t}$ are predefined noise schedules ensuring $q_{T}(x_{T}|x_{0})\approx q_{T}(x_{T})\approx\mathcal{N}(x_{T}|0,\tilde{\sigma}^{2}I)$ for some $\tilde{\sigma}>0$ that is independent of $x_{0}$.
  • Backward process. Starting from $x_{T}\sim\mathcal{N}(x_{T}|0,\tilde{\sigma}^{2}I)$, diffusion models reconstruct the original data $x_{0}$ by solving diffusion ODE/SDE [49] from $T$ to 0:
    $\text{(Diffusion ODE) }\quad dx_{t}=\left[f(t)x_{t}-\frac{1}{2}g^{2}(t)\nabla_{x_{t}}\log q_{t}(x_{t})\right]dt,$ (20)
    $\text{(Diffusion SDE) }\quad dx_{t}=\left[f(t)x_{t}-g^{2}(t)\nabla_{x_{t}}\log q_{t}(x_{t})\right]dt+g(t)d\tilde{\omega}$, (21)
    where $f(t)=\frac{d\log\alpha_{t}}{dt}$, $g^{2}(t)=\frac{d\sigma_{t}^{2}}{dt}-2\frac{d\log\alpha_{t}}{dt}\sigma_{t}^{2}$ [28], and $\tilde{\omega}$ is a standard Wiener process when time flows backwards from $T$ to 0.
Diffusion models train a neural network $\epsilon_{\theta}(x_{t},t)$ parameterized by $\theta$ with the following objective [22, 37]:
$\min_{\theta}\mathbb{E}_{x_{t},\,t\sim\mathcal{U}([0,T]),\epsilon\sim\mathcal{N}(0,I)}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\right],$ (22)
where $x_{0}\sim q_{0}(x_{0})$, $x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon$. The fitted $\theta$ can be use to estimate the score function $\nabla_{x_{t}}\log q_{t}(x_{t})$ by $\epsilon_{\theta}(x_{t},t)\approx-\sigma_{t}\nabla_{x_{t}}\log q_{t}(x_{t})$ and substituted into Eq. (20-21) for $x_{0}$ generation.
Energy-guided diffusion sampling. To bias sampling toward preferred distributions, the prior energy-guidance framework [39] provides the re-weighted distribution $p_{0}(x_{0})$:
$p_{0}(x_{0})\propto q_{0}(x_{0})f(\mathcal{E}(x_{0})),$ (23)
where $q_{0}(x_{0})$ is the unknown data distribution, $\mathcal{E}(x_{0})$ is any form of energy function that encodes human preferences, and $f(x)\geq 0$ can be any non-negative function. To solve this sampling problem, previous approaches train a separate time-dependent classifier $\mathcal{E}_{t}$ [39, 24], which is computationally expensive and may introduce additional training errors. To address this, previous work [72] introduces weighted regression method to perform exact energy guidance:
Lemma 1 (Weighted regression as exact energy guidance).
We can sample $x_{0}\sim p_{0}(x_{0})$ where $p_{0}$ is the weighted distribution in Eq. (23) by (i) optimizing the weighted regression loss:
$\min_{\theta}\mathbb{E}_{x_{t},t\sim\mathcal{U}([0,T]),\epsilon\sim\mathcal{N}(0,I)}\left[f\big(\mathcal{E}(x_{0})\big)\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\right],$ (24)
where $x_{0}\sim q_{0}(x_{0})$, $x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon$; and (ii) substituting the obtained $\epsilon_{\theta}$ into diffusion ODEs/SDEs [49] solving process.
Task-conditioned diffusion sampling. In this work, we extend the Lemma 1 and provide the following proposition for task-specific sampling:
Proposition 2 (Task-conditioned diffusion policy extraction via weighted regression).
The extraction of optimal policy $\pi_{z}^{*}$ in Eq. (11) can be achieved by (i) minimizing the weighted regression loss defined as:
$\min_{\theta}\mathbb{E}_{t\sim\mathcal{U}([0,T]),\epsilon\sim\mathcal{N}(0,I),(s,a)\sim\mathcal{D}}\Big[\exp\Big(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\Big)\|\epsilon-\epsilon_{\theta,z}(a_{t},s,z,t)\|_{2}^{2}\Big],$
and (ii) solving diffusion ODEs/SDEs [49] by substituting $\epsilon_{\theta,z}$ in Eq. (20-21).
Proof.
We aim to extract the optimal policy forming as a weighted distribution as follows:
$\pi_{z}^{*}(a|s)\propto\mu(a|s)\exp\left(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\right),$
where $\mu(a|s)$ is the behavior distribution in dataset $\mathcal{D}$ and $F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)$ is the task-dependent weight function. Set $p_{z}(a|s)$ as a task-conditioned reweighted version of the unknown data distribution $q_{0}(a|s)$ and substitute the task-dependent weight function into Eq. (23), we have:
$p_{z}(a|s)\propto q_{0}(a|s)\exp\left(\alpha\cdot\left(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)\right)\right).$ (25)
There exist a normalization constant $Z(s,z)=\int q_{0}(a|s)\exp\left(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\right)da$ ensures $p_{z}(a|s)$ to be a valid distribution. We now aim to sample from the distribution $p_{z}(a|s)$.
Parameterize the denoising process with a task-conditioned model $\epsilon_{\theta,z}(a_{t},s,z,t)$ and substitute in Eq. (24), we can get a modified task-conditioned weighted regression objective for the diffusion models by replacing the energy function with $F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)$:
$\min_{\theta}\mathbb{E}_{t\sim\mathcal{U}([0,T]),\epsilon\sim\mathcal{N}(0,I),(s,a)\sim\mathcal{D}}\Big[\exp\Big(\alpha\cdot(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z))\Big)\|\epsilon-\epsilon_{\theta,z}(a_{t},s,z,t)\|_{2}^{2}\Big].$ (26)
According to the Lemma 1, substituting $\epsilon_{\theta,z}^{*}$ into the diffusion ODEs/SDEs [49] generates data that sample from $p_{z}(a|s)$, thereby recovering $\pi_{z}^{*}(a|s)$. The proof is completed. ∎
Appendix B Experimental Setups
Our experiments span two benchmarks, ExORL [70] and D4RL [14], across two locomotion domains (Walker and Quadruped) and two manipulation goal-reaching domains (Jaco and Kitchen). All experiments we conducted are in state-based tasks.
B.1 Datasets
ExORL [70].
ExORL consists of datasets collected by several unsupervised RL algorithms [34] on the DeepMind Control Suite [54]. We select datasets collected by four unsupervised RL algorithms: APS [38], RND [5], PROTO [69], and DIAYN [11] for each of the three domains (Walker, Jaco, and Quadruped). Each domain has four tasks for evaluation at test time.
D4RL [14] Kitchen.
D4RL Kitchen datasets consist of robotic arm manipulating trajectories of different tasks in the Franka Kitchen [19] domain. We adopt two datasets: "kitchen-partial-v0" and "kitchen-mixed-v0".
  • Partial: Subtasks involve the microwave, kettle, light switch, and slide cabinet. Tasks are guaranteed to be solved in a subset of the ’partial’ dataset.
  • Mixed: Subtasks involve microwave, kettle, bottom burner, and light switch. no trajectories solve any tasks completely in the ’mixed’ dataset.
Dataset preprocessing.
BREEZE incorporates standardized data normalization for diffusion models across all training datasets as a preprocessing step to stabilize the learning process and enhance generalization capability.
B.2 Domains
Figure 8: Domains. Walker, Quadruped, Jaco, and Franka Kitchen.
  • Walker (Locomotion): A bipedal robot with 24-dimensional states (joint positions/velocities) and 6-dimensional actions. Test tasks include Flip, Run, Stand, and Walk. Rewards combine dense objectives: maintaining torso height (Stand), achieving target velocities (Run/Walk), or angular momentum (Flip).
  • Quadruped (Locomotion): A four-legged robot with 78-dimensional states and 12-dimensional actions. Tasks include Jump, Run, Stand, and Walk, with rewards for torso stability and velocity tracking.
  • Jaco (Manipulation/Goal-reaching): A 6-DoF robotic arm with 55-dimensional states and 6-dimensional actions. Tasks involve reaching four target positions (Top/Bottom Left/Right) using sparse rewards based on proximity to goals.
  • Franka Kitchen [19] (Manipulation/Goal-reaching): A robotic manipulation environment controlling a 9-DoF Franka robot with multiple given test-time objectives. The agent must sequentially achieve four subtasks per episode, receiving sparse rewards for each subtask. States include arm joint positions, velocities, torques, and task-specific object features. Test-time goals are defined as proprioceptive states concatenated with target object states.
B.3 Baselines
We utilize the following open-source codebases for experiments on baselines. We set the batch size to 512 to ensure a fair comparison with other methods, while keeping all other hyperparameters at their default values.
B.4 Architectures
This section details the model architectures used in BREEZE.
$F$ Network.
The $F$ network processes inputs $(s,a)$ and $(s,z)$ using two separate MLPs, each with two hidden layers of size 1024 and ReLU activation, projecting the inputs into a 512-dimensional feature space. The resulting embeddings are concatenated along the sequence dimension and processed through two identical processing blocks. Each block consists of a self-attention layer [59], followed by a feed-forward network with residual connections, LayerNorm [2], and dropout (rate=0.1). Specifically, the final representation is flattened and passed through two independent output heads, denoted as $F1$ and $F2$, each comprising two linear layers with a hidden dimension of 256 that map to the $d$-dimensional output space.
$B$ Network.
The $B$ network is based on a standard transformer architecture with 8-head multi-head attention [59] and ReLU activation. State embeddings are processed through transformer blocks and projected to a $d$-dimensional space via a linear layer and further scaled to $\sqrt{d}$ by L2 normalization. We observe that a sufficiently large $B$ network is essential for achieving strong performance on the ExORL benchmark.
Diffusion Policy.
The diffusion policy uses a residual-connected MLP as the noise predictor, following the IDQL [20] implementation. The hidden dimension is set to 1024. The task vector $z$ is incorporated via Feature-wise Linear Modulation (FiLM) [46]: two parallel linear layers generate a scaling factor $\gamma$ and a shifting factor $\beta$ from $z$, which modulate intermediate features $h$ as $\gamma\odot h+\beta$.
B.5 Sampling of $z$
Following previous studies [58, 26], we utilize a mixture of two methods for sampling the latent variable $z$:
  • Sample $z$ uniformly within a sphere of radius $\sqrt{d}$ in $\mathbb{R}^{d}$.
  • Derive $z$ by setting $z=B(s)$, where $s\sim\mathcal{D}$ is randomly sampled from the data set.
In our primary experiments, we employ a balanced mix ratio of 50% between the two methods.
B.6 Value Learning
Training details.
We implement the $V$-network with 3-layer MLPs with 512 hidden dimensions and ReLU activation functions. During training, we set the $\tau$ for expectile regression in Eq. (7) shown in Section  B.7. We use clipped double value learning [16] for both the $Q$-value and the $M$-value.
Rejection sampling.
To improve the zero-shot performance, we use rejection sampling rather than tuning existing parameters for greedy optimization. Rejection sampling is commonly used in diffusion policies [7, 20, 21], a stable mechanism that selects $N$ actions from the policy to boost with the highest value. Our practical implementation of rejection sampling is in two phases:
  • Training phase: To compute the FB loss (Eq. 4), we first sample a full transition batch, which contains states, actions, next states, and next actions. A hyperparameter $\rho_{a}$, referred to as the dataset mixture ratio, determines the proportion of next actions drawn directly from the dataset. The remaining fraction is replaced by actions sampled from the policy $\pi_{z}$: we generate $K_{\text{train}}$ candidate actions $a_{t+1}^{(1)},\dots,a_{t+1}^{(K_{\text{train}})}$} via the diffusion policy and select the optimal next action according to:
    $a_{t+1}^{*}=\arg\max_{a_{t+1}^{(i)}};F(s_{t+1},a_{t+1}^{(i)},z)^{\top}z.$ (27)
  • Evaluation phase: We sample $K_{\text{eval}}$ candidate actions during evaluation, and select the optimal action according to the $Q$-value.
Details of practical $\rho_{a}$, $K_{\text{train}}$ and $K_{\text{eval}}$ is provided in Section B.7.
B.7 Hyperparameter
This section provides the detailed hyperparameter setup. In our experiments, the model architecture and basic algorithm hyperparameters remain unchanged, as detailed in Table 5. Domain-specific hyperparameters are detailed in Table 6 and Table 7.
Table 5: General hyperparameters used for BREEZE
General hyperparameters Hyperparameter Value
Optimizer AdamW
Representation learning rate 1e-4
$V$ network learning rate 3e-4
Policy learning rate 1e-4
Discount factor $\gamma$ 0.98
Learning steps 1,000,000
Mini-batch 512
Representation soft update factor $\lambda$ 0.01
Policy soft update factor $\lambda$ 0.001
Latent dimension $d$ 50
$z$ mixing ratio 0.5
$z$ inference steps 10,000
Temperature $\alpha$ 0.05 (0.1 kitchen-mixed)
Regularization weight for orthogonality loss $w_{1}$ 10
Beta schedule vp
Architecture $F$ preprocessor hidden dimension 1024
$F$ preprocessor hidden layers 2
$F$ preprocessor output dimension 512
$F$ preprocessor activation function relu
$F$ attention blocks 2
$F$ linear layer hidden dimension 256
$F$ dropout 0.1
$B$ nhead 8
$B$ encoder layers 2
$B$ decoder layers 2
$B$ d model 256
$B$ dropout 0.1
$B$ feedforward dimension 2048
$B$ activation function relu
$V$ hidden dimension 512
$V$ hidden layers 2
$V$ activation function relu
Policy MLP blocks 3
Policy hidden dimension 1024
Policy activation function mish
Policy time embedding learned
Table 6: Domain-specific hyperparameters for BREEZE with full dataset.
Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$
Walker-RND 9 0.1 64 0.99 0.001 5
Walker-APS 9 0.1 64 0.99 0.001 5
Walker-PROTO 9 0.1 64 0.99 0.001 5
Walker-DIAYN 9 0.1 64 0.99 0.001 5
Jaco-RND 1 0 64 0.99 0.001 10
Jaco-APS 9 0.1 64 0.99 0.0001 5
Jaco-PROTO 1 0 64 0.99 0.001 10
Jaco-DIAYN 9 0.1 64 0.99 0.0001 5
Quadruped-RND 1 0 64 0.99 0.001 10
Quadruped-APS 1 0 64 0.99 0.001 10
Quadruped-PROTO 1 0 64 0.99 0.001 10
Quadruped-DIAYN 1 0 64 0.99 0.001 10
Kitchen-mixed 1 0 16 0.7 0.001 5
Kitchen-partial 2 0.2 8 0.7 0.001 5
Table 7: Domain-specific hyperparameters for BREEZE with small sample dataset.
Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$
Walker-RND 2 0.2 32 0.99 0.0001 10
Walker-APS 2 0.2 32 0.99 0.0001 10
Walker-PROTO 2 0.2 32 0.99 0.001 10
Walker-DIAYN 8 0.2 32 0.99 0.001 10
Jaco-RND 1 0 32 0.99 0.0001 10
Jaco-APS 1 0 32 0.99 0.0001 10
Jaco-PROTO 1 0 32 0.99 0.0001 10
Jaco-DIAYN 1 0 32 0.99 0.0001 10
Quadruped-RND 2 0.2 32 0.99 0.001 10
Quadruped-APS 1 0 32 0.99 0.001 10
Quadruped-PROTO 2 0.2 32 0.99 0.001 10
Quadruped-DIAYN 2 0.2 32 0.99 0.0001 10
Appendix C Pseudo-Code
Algorithm 1 provides the pseudocode. Our implementation uses PyTorch [45], with all experiments conducted on a single NVIDIA A6000 GPU. Peak memory usage under 23 GB.
Algorithm 1 Behavior-Regularized Zero-shot RL (BREEZE)
1:Latent dimension $d$, mini-batch $b$, gradient steps NN, orthogonality coefficient $w_{1}$, value regularization coefficient $w_{q}$.
2:Normalize dataset $\mathcal{D}$; Initialize networks $F_{\phi}$, $B_{\psi}$, $V_{\omega}$, diffusion actor $\pi_{z}$ with network $\epsilon_{\theta}$
3:for $n=1,\ldots,N$ do
4:  Sample a mini-batch of transitions $\{(s_{t},a_{t},s_{t+1},a_{t+1})_{i}\}_{i\in|b|}\subset\mathcal{D}$
5:  Sample a mini-batch of $\{z_{i}\}_{i\in|b|}\sim\mathbb{R}^{d}$
6:  Collect $\{a^{\prime}_{t+1}\}_{i\in|b|}$ by mixture of data batch and $\pi_{z}$ outputs
7:  // Representation learning:
8:  Compute $\mathcal{L}_{\text{FB}}(\phi,\psi)$ and $\mathcal{L}_{F\text{-reg}}(\phi)$ using Eq. (4), Eq. (9)
9:  Compute orthogonality regularization loss $\mathcal{L}_{\text{ortho}}$
10:  $\mathcal{L}_{\mathrm{ortho}}(\psi)=\frac{1}{|b|^{2}}\sum_{i,j\in|b|}\left[(B_{\psi}(s_{i})^{\top}B_{\psi}(s_{j}^{\prime}))^{2}-\|B_{\psi}(s_{i})\|_{2}^{2}-\|B_{\psi}(s_{j}^{\prime})\|_{2}^{2}\right]$
11:  Update $F_{\phi}$ and $B_{\psi}$ by $\mathcal{L}_{\text{FB}}+w_{1}\cdot\mathcal{L}_{\text{ortho}}+w_{q}\cdot\mathcal{L}_{F\text{-reg}}$
12:  // Value function learning:
13:  Update $V_{\omega}$ using Eq. (7)
14:  // Diffusion policy learning:
15:  Update ${\epsilon}_{\theta}$ using Eq. (12)
16:end for
Appendix D Additional Results
This section provides a comprehensive showcase of all experimental results.
  • Section D.1 provides the learning curve comparisons.
  • Section D.2 reports the detailed results on ExORL benchmark.
  • Section D.3 presents the visualizations of our empirically evaluated value distributions.
D.1 Learning Curves
This section presents the zero-shot learning curves on the ExORL benchmark. Results on the full dataset and the 100k small-sample dataset are shown in Figure 9-11 and Figure 12-14, respectively. The solid lines represent the mean IQM returns over 5 random seeds, and the shaded regions correspond to the standard deviation.
Figure 9: Curves of zero-shot performance on Quadruped domain.
Figure 10: Curves of zero-shot performance on Jaco domain.
Figure 11: Curves of zero-shot performance on Walker domain.
Figure 12: Curves of zero-shot performance on Quadruped domain with 100k dataset.
Figure 13: Curves of zero-shot performance on Jaco domain with 100k dataset.
Figure 14: Curves of zero-shot performance on Walker domain with 100k dataset.
Figure 15: Curves of zero-shot performance on Kitchen domain.
D.2 Detailed Results on ExORL Benchmarks
This section presents complete experimental results on the ExORL benchmark. Table 8 summarizes performance on the full dataset, while Table 9 reports results on the 100k small-sample datasets. Evaluations were conducted every 10,000 steps during offline training. Following standard zero-shot RL criteria, we computed the aggregate score across all tasks as the overall performance measure, and selected the highest domain-level score across 5 random seeds along with corresponding per-task results.
Table 8: Full dataset experimental results on ExORL
Dataset Domain Task SF-LAP FB VCFB MCFB HILP BREEZE
RND Walker walk $555\pm 221$ $918\pm 26$ $813\pm 55$ $842\pm 126$ $798\pm 29$ $907\pm 23$
stand $810\pm 47$ $834\pm 39$ $913\pm 29$ $950\pm 20$ $888\pm 55$ $938\pm 12$
run $235\pm 61$ $357\pm 20$ $376\pm 22$ $328\pm 20$ $368\pm 34$ $320\pm 25$
flip $465\pm 107$ $535\pm 29$ $511\pm 52$ $517\pm 89$ $608\pm 48$ $606\pm 22$
whole performance $516\pm 97$ $661\pm 10$ $653\pm 22$ $59\pm 51 66$$65\pm 33$ $693\pm 16$
Jaco reach top right $2\pm 2$ $75\pm 68$ $78\pm 63$ $86\pm 68$ $43\pm 42$ $108\pm 42$
reach top left $2\pm 2$ $4\pm 8$ $15\pm 16$ $10\pm 9$ $53\pm 17$ $68\pm 30$
reach bottom right $1\pm 0$ $46\pm 42$ $81\pm 62$ $56\pm 56$ $53\pm 41$ $77\pm 45$
reach bottom left $70\pm 71$ $4\pm 5$ $9\pm 10$ $11\pm 13$ $60\pm 39$ $84\pm 55$
whole performance $18\pm 18$ $32\pm 23$ $46\pm 35$ $41\pm 34$ $52\pm 21$ $84\pm 14$
Quadruped walk $206\pm 136$ $519\pm 35$ $462\pm 6$ $603\pm 68$ $629\pm 69$ $641\pm 64$
stand $539\pm 224$ $956\pm 11$ $883\pm 52$ $965\pm 5$ $904\pm 41$ $936\pm 7$
run $233\pm 136$ $484\pm 14$ $431\pm 20$ $486\pm 22$ $484\pm 20$ $514\pm 22$
jump $342\pm 238$ $722\pm 4$ $659\pm 53$ $683\pm 41$ $679\pm 72$ $808\pm 16$
whole performance $330\pm 165$ $671\pm 14$ $609\pm 26$ $684\pm 18$ $674\pm 28$ $725\pm 23$
APS Walker walk $272\pm 81$ $725\pm 118$ $705\pm 130$ $817\pm 88$ $802\pm 22$ $935\pm 23$
stand $663\pm 54$ $720\pm 208$ $567\pm 109$ $751\pm 65$ $860\pm 10$ $865\pm 57$
run $139\pm 25$ $244\pm 58$ $245\pm 53$ $16\pm 26 33$$32\pm 26$ $267\pm 14$
flip $221\pm 35$ $380\pm 54$ $431\pm 78$ $430\pm 97$ $578\pm 49$ $482\pm 52$
whole performance $324\pm 24$ $517\pm 99$ $487\pm 75$ $578\pm 35$ $643\pm 22$ $637\pm 21$
Jaco reach top right $29\pm 22$ $33\pm 37$ $18\pm 10$ $28\pm 8$ $98\pm 44$ $119\pm 52$
reach top left $9\pm 15$ $6\pm 14 11$$1\pm 11$ $15\pm 17$ $80\pm 28$ $153\pm 76$
reach bottom right $50\pm 32$ $33\pm 27$ $37\pm 46$ $31\pm 21$ $105\pm 26$ $70\pm 34$
reach bottom left $70\pm 72$ $6\pm 8$ $13\pm 11$ $15\pm 8$ $52\pm 20$ $186\pm 26$
whole performance $39\pm 26$ $22\pm 14$ $0\pm 18 22$$2\pm 3$ $84\pm 16$ $132\pm 16$
Quadruped walk $312\pm 158$ $615\pm 48$ $541\pm 38$ $601\pm 90$ $557\pm 40$ $660\pm 92$
stand $796\pm 186$ $898\pm 57$ $946\pm 24$ $895\pm 63$ $934\pm 9$ $950\pm 12$
run $342\pm 199$ $471\pm 32$ $76\pm 10 44$$40\pm 36$ $466\pm 13$ $472\pm 27$
jump $542\pm 126$ $686\pm 33$ $692\pm 6$ $699\pm 70$ $758\pm 21$ $710\pm 27$
whole performance $498\pm 160$ $68\pm 29 66$$64\pm 3$ $659\pm 50$ $679\pm 14$ $698\pm 24$
PROTO Walker walk $352\pm 210$ $889\pm 45$ $783\pm 165$ $867\pm 109$ $894\pm 38$ $923\pm 2$
stand $703\pm 155$ $936\pm 30$ $832\pm 86$ $845\pm 135$ $931\pm 22$ $858\pm 35$
run $202\pm 69$ $298\pm 27$ $324\pm 81$ $262\pm 69$ $387\pm 29$ $228\pm 26$
flip $271\pm 156$ $477\pm 47$ $503\pm 54$ $435\pm 143$ $647\pm 76$ $645\pm 62$
whole performance $382\pm 129$ $650\pm 19$ $611\pm 94$ $602\pm 112$ $715\pm 31$ $663\pm 19$
Jaco reach top right $5\pm 9$ $31\pm 54$ $12\pm 17$ $37\pm 52$ $55\pm 23$ $80\pm 75$
reach top left $2\pm 2$ $33\pm 34$ $7\pm 9$ $14\pm 13$ $52\pm 28$ $99\pm 48$
reach bottom right $29\pm 46$ $19\pm 33$ $25\pm 36$ $26\pm 40$ $30\pm 24$ $91\pm 79$
reach bottom left $24\pm 41$ $0\pm 0$ $8\pm 13$ $2\pm 2$ $38\pm 33$ $26\pm 29$
whole performance $15\pm 14$ $21\pm 26$ $13\pm 12$ $20\pm 21$ $44\pm 19$ $74\pm 26$
Quadruped walk $131\pm 32$ $168\pm 76$ $136\pm 63$ $235\pm 144$ $118\pm 58$ $298\pm 79$
stand $293\pm 88$ $351\pm 263$ $239\pm 95$ $259\pm 132$ $329\pm 95$ $569\pm 55$
run $163\pm 51$ $125\pm 112$ $100\pm 45$ $213\pm 172$ $175\pm 44$ $283\pm 33$
jump $211\pm 77$ $244\pm 92$ $265\pm 166$ $169\pm 101$ $242\pm 70$ $407\pm 7$
whole performance $199\pm 10$ $222\pm 107$ $185\pm 72$ $219\pm 135$ $216\pm 54$ $389\pm 44$
DIAYN Walker walk $174\pm 151$ $364\pm 164$ $329\pm 147$ $348\pm 174$ $604\pm 126$ $613\pm 103$
stand $558\pm 129$ $614\pm 146$ $484\pm 194$ $429\pm 83$ $696\pm 168$ $712\pm 66$
run $96\pm 25$ $103\pm 20$ $74\pm 26$ $75\pm 25$ $156\pm 25$ $152\pm 38$
flip $128\pm 24$ $272\pm 58$ $183\pm 51$ $222\pm 112$ $386\pm 76$ $373\pm 39$
whole performance $239\pm 79$ $338\pm 74$ $268\pm 67$ $268\pm 97$ $461\pm 64$ $463\pm 42$
Jaco reach top right $43\pm 37$ $17\pm 12$ $24\pm 16$ $1\pm 1$ $64\pm 5$ $76\pm 14$
reach top left $10\pm 12$ $18\pm 7$ $30\pm 17$ $5\pm 2 55$$5\pm 6$ $58\pm 33$
reach bottom right $42\pm 42$ $27\pm 17$ $7\pm 4$ $1\pm 1$ $45\pm 26$ $96\pm 36$
reach bottom left $32\pm 41$ $28\pm 12$ $36\pm 12$ $4\pm 3$ $46\pm 4$ $82\pm 35$
whole performance $32\pm 26$ $22\pm 6$ $24\pm 3$ $15\pm 1$ $52\pm 7$ $78\pm 11$
Quadruped walk $196\pm 151$ $475\pm 77$ $311\pm 47$ $596\pm 26$ $497\pm 18$ $482\pm 11$
stand $305\pm 270$ $756\pm 30$ $793\pm 87$ $865\pm 30$ $4\pm 4 944±9$$4\pm 9$
run $139\pm 121$ $380\pm 22$ $364\pm 22$ $408\pm 15$ $469\pm 2$ $472\pm 5$
jump $190\pm 158$ $636\pm 19$ $576\pm 47$ $700\pm 18$ $769\pm 8$ $766\pm 25$
whole performance $207\pm 168$ $562\pm 23$ $511\pm 37$ $643\pm 14$ $70\pm 4 66$$66\pm 2$
Table 9: Small sample dataset experimental results on ExORL.
Dataset Domain Task FB VCFB MCFB BREEZE (ours)
RND Walker walk $300\pm 50$ $377\pm 18$ $225\pm 46$ $663\pm 41$
stand $368\pm 70$ $556\pm 44$ $587\pm 97$ $791\pm 21$
run $104\pm 11$ $144\pm 17$ $101\pm 18$ $224\pm 12$
flip $284\pm 91$ $319\pm 46$ $236\pm 44$ $421\pm 9$
whole performance $264\pm 33$ $350\pm 29$ $287\pm 48$ $525\pm 13$
Jaco reach top right $9\pm 8$ $6\pm 7$ $28\pm 16$ $51\pm 10$
reach top left $2\pm 2$ $6\pm 5$ $1\pm 1$ $7\pm 3$
reach bottom right $10\pm 10$ $4\pm 4$ $19\pm 15$ $38\pm 22$
reach bottom left $8\pm 9$ $20\pm 12$ $1\pm 1$ $47\pm 22$
whole performance $7\pm 5$ $9\pm 2$ $13\pm 7$ $36\pm 5$
Quadruped walk $125\pm 98$ $164\pm 45$ $94\pm 58$ $382\pm 15$
stand $268\pm 183$ $288\pm 73$ $159\pm 70$ $651\pm 37$
run $132\pm 83$ $186\pm 52$ $89\pm 49$ $318\pm 16$
jump $178\pm 130$ $292\pm 82$ $148\pm 70$ $544\pm 37$
whole performance $176\pm 123$ $233\pm 52$ $123\pm 61$ $474\pm 21$
APS Walker walk $535\pm 139$ $533\pm 114$ $449\pm 64$ $762\pm 25$
stand $447\pm 45$ $625\pm 52$ $562\pm 104$ $828\pm 19$
run $166\pm 73$ $216\pm 33$ $192\pm 50$ $200\pm 2$
flip $334\pm 121$ $293\pm 85$ $354\pm 95$ $365\pm 24$
whole performance $370\pm 66$ $416\pm 10$ $389\pm 77$ $539\pm 15$
Jaco reach top right $36\pm 29$ $7\pm 7$ $20\pm 24$ $21\pm 16$
reach top left $3\pm 2$ $7\pm 5$ $24\pm 19$ $21\pm 15$
reach bottom right $44\pm 60$ $42\pm 43$ $56\pm 71$ $58\pm 33$
reach bottom left $2\pm 3$ $0\pm 0$ $15\pm 4$ $51\pm 29$
whole performance $21\pm 17$ $14\pm 13$ $29\pm 27$ $38\pm 9$
Quadruped walk $299\pm 27$ $293\pm 26$ $255\pm 116$ $435\pm 49$
stand $479\pm 12$ $524\pm 141$ $456\pm 92$ $810\pm 85$
run $260\pm 10$ $32\pm 46 22$$22\pm 101$ $401\pm 30$
jump $321\pm 115$ $56\pm 30 33$$38\pm 179$ $577\pm 59$
whole performance $340\pm 29$ $351\pm 57$ $318\pm 122$ $556\pm 52$
PROTO Walker walk $522\pm 136$ $732\pm 38$ $685\pm 87$ $709\pm 27$
stand $608\pm 119$ $709\pm 36$ $637\pm 39$ $846\pm 22$
run $176\pm 58$ $215\pm 24$ $174\pm 49$ $210\pm 7$
flip $354\pm 67$ $394\pm 36$ $357\pm 43$ $447\pm 27$
whole performance $415\pm 19$ $513\pm 31$ $463\pm 11$ $553\pm 18$
Jaco reach top right $7\pm 10$ $17\pm 19$ $6\pm 8$ $18\pm 13$
reach top left $59\pm 4$ $14\pm 12$ $19\pm 14$ $51\pm 22$
reach bottom right $0\pm 0$ $22\pm 22$ $12\pm 10$ $27\pm 19$
reach bottom left $0\pm 0$ $18\pm 11$ $12\pm 10$ $19\pm 12$
whole performance $16\pm 2$ $18\pm 12$ $12\pm 7$ $29\pm 12$
Quadruped walk $105\pm 63$ $22\pm 7$ $103\pm 85$ $129\pm 66$
stand $327\pm 180$ $196\pm 202$ $387\pm 185$ $59\pm 35$
run $169\pm 99$ $98\pm 110$ $206\pm 152$ $244\pm 173$
jump $193\pm 111$ $106\pm 103$ $265\pm 194$ $291\pm 47$
whole performance $198\pm 111$ $106\pm 103$ $240\pm 134$ $181\pm 60$
DIAYN Walker walk $03\pm 79 11$$14\pm 16$ $129\pm 38$ $418\pm 119$
stand $513\pm 245$ $460\pm 256$ $424\pm 27$ $524\pm 39$
run $87\pm 45$ $79\pm 22$ $63\pm 19$ $102\pm 17$
flip $106\pm 50$ $186\pm 44$ $169\pm 33$ $276\pm 28$
whole performance $202\pm 94$ $210\pm 81$ $196\pm 26$ $330\pm 43$
Jaco reach top right $18\pm 23$ $14\pm 13$ $29\pm 37$ $5\pm 34$
reach top left $32\pm 35$ $25\pm 9$ $9\pm 8$ $72\pm 58$
reach bottom right $15\pm 11$ $12\pm 9$ $42\pm 58$ $6\pm 29$
reach bottom left $5\pm 2$ $19\pm 24$ $0\pm 0$ $7\pm 26$
whole performance $17\pm 11$ $18\pm 5$ $0\pm 26 22$$2\pm 15$
Quadruped walk $158\pm 41$ $189\pm 73$ $190\pm 10$ $303\pm 69$
stand $430\pm 84$ $423\pm 46$ $408\pm 57$ $680\pm 117$
run $224\pm 41$ $214\pm 28$ $204\pm 28$ $341\pm 79$
jump $368\pm 83$ $327\pm 58$ $343\pm 63$ $459\pm 61$
whole performance $295\pm 46$ $288\pm 48$ $286\pm 34$ $446\pm 78$
D.3 Empirically Evaluated $M^{\pi_{z}}$ and $Q_{z}$ on Other Tasks
We present empirical visualization results of $M^{\pi_{z}}$ and $Q_{z}$ distributions, shown in Figure 18-18. We trained FB, VCFB, and MCFB on the RND [5] dataset. Specifically, the trained $B$ network was utilized for inferring the task latent vector $z_{\text{task}}$, and as the input to both $M^{\pi_{z}}$ and $Q_{z}$ during the evaluation phase.
Figure 16: Visualization of the empirical $M^{\pi_{z}}$ and $Q_{z}$ distribution on Walker RND.
Figure 17: Visualization of the empirical $M^{\pi_{z}}$ and $Q_{z}$ distribution on Jaco RND.
Figure 18: Visualization of the empirical $M^{\pi_{z}}$ and $Q_{z}$ distribution on Quadruped RND.
Appendix E Discussion and Limitation
E1. Hyperparameter Sensitivity and Tuning.
BREEZE introduces additional components compared to vanilla FB, resulting in more hyperparameters. While performance is generally robust, we provide the following tuning guidance:
  • Effect of components.
    Representation architectures are responsible for a high performance ceiling, as we can observe from Table 4. The diffusion policy ensures a solid performance lower bound by maintaining stable alignment with any quality dataset.
  • Greedy demand for upstream value learning.
    Training cost is mainly influenced by the action ratio $\rho_{a}$ and the number of rejection sampling candidates $K_{\text{train}}$. While the diffusion policy provides conservatism, we find that increasing $\rho_{a}$ and $K_{\text{train}}$ further improves performance in some domains at the cost of additional computation. Near-optimal results can be achieved cost-effectively by reducing the batch size and diffusion timesteps to reallocate the computational resources to increase $\rho_{a}$ and $K_{\text{train}}$.
E2. Computation and Performance Trade-off.
A limitation of BREEZE is its higher computational cost, attributable to the expressive representation networks and diffusion model. We consider this a reasonable trade-off given the significant improvements in stability and zero-shot performance. For a comparable and fair comparison, all reported results use a batch size of 512, and training for 1 million steps takes approximately 20 hours or less for environments with standard TD updates [50]; configurations with higher $K_{\text{train}}$ and mixture ratios require up to 39 hours. As shown in Table 10, BREEZE typically converges within 400k steps, matching or exceeding baselines trained for 1M steps. Future work may explore framework-level optimizations, such as replacing diffusion with more efficient flow-based policies to lower costs without compromising performance.
Table 10: Computation cost v.s. performance trade-off. Aggregated scores are averaged over 5 random seeds across datasets. Experiments are conducted on a single NVIDIA A6000 GPU.
Methods Training Steps (k) Training Time (h) IQM Walker IQM Jaco IQM Quadruped Aggregate IQM
[…]
400 1.6 508 21 401 310
1000 4.0 542 24 531 366
[…]
400 4.8 489 19 457 322
1000 12.0 505 26 492 341
[…]
400 4.8 514 21 469 335
1000 12.0 527 25 551 368
[…]
400 8.0-15.6 606 75 567 416

### v2

@@
 However, its real-world adoption faces two persistent challenges: the reliance on human-provided reward functions and its task-specific learning paradigm, which limits adaptability to novel or multiple tasks.
-These challenges sparked growing interests in zero-shot RL [57, 58, 44, 26, 4, 23, 56, 13], which enables learning a versatile agent through pretraining on reward-free transitions, and then zero-shot adapting to arbitrary reward functions during inference.
+These challenges sparked growing interest in zero-shot RL [57, 58, 44, 26, 4, 23, 56, 13], which enables learning a versatile agent through pretraining on reward-free transitions and then zero-shot adaptation to arbitrary reward functions during inference.
 This opens up new possibilities for developing general-purpose RL agents capable of generalizing across diverse tasks in open-world scenarios.
@@
 This preserves the task generalization capability of policies, but often results in heavy, manually designed pretraining without optimality guarantees.
-The second category of methods adopt a more principled approach by decomposing the problem into dynamic representations that can be recomposed for novel tasks without retraining.
+The second category of methods adopts a more principled approach by decomposing the problem into dynamic representations that can be recomposed for novel tasks without retraining.
 Among dynamic representation-based methods, Forward-Backward representations (FB) [57, 58] have recently attracted notable attention, factorizing occupancy measures into two components: a forward representation that captures policy dynamics and a backward representation that encodes global state information.
@@
 First, learning complex approximators and multimodal behaviors requires highly expressive models, which current FB methods lack in both their representations and policy.
-Second, the offline, unsupervised pretraining stage suffers from extrapolation error due to out-of-distribution (OOD) actions, which is similar to offline RL [17, 32, 33], but exhibits more complex behavior.
+Second, the offline, unsupervised pretraining stage suffers from extrapolation error due to out-of-distribution (OOD) actions, which is a similar problem to offline RL [17, 32, 33], but exhibits more complex behavior.
 As our results show, naively integrating value constraints can be ineffective [26], indicating a need for a more delicate OOD regularization mechanism.
@@
 Finally, we employ expressive representation networks based on the attention architecture to capture complicated dynamics.
-We conducted extensive experiments on the ExORL [70] benchmark and the D4RL [14] Kitchen dataset, under both full datasets and small-sample data regimes.
+We conducted extensive experiments on the ExORL benchmark [70] and the D4RL Kitchen dataset [14], under both full datasets and small-sample data regimes.
 The results demonstrate that BREEZE achieves the best or near-the-best performance while exhibiting superior robustness.
@@
 The $Q$-function and $V$-function for $\pi$ starting at $s_{0},a_{0}$ under a given reward function $r:\mathcal{S}\to\mathbb{R}$ are respectively defined as $Q^{\pi}_{r}(s_{0},a_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},a_{0},\pi]$ and $V^{\pi}_{r}(s_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},\pi]$.
-The goal of the zero-shot RL problem is to train a task-agnostic policy given offline dataset $\mathcal{D}=\{s_{i},a_{i},s_{i+1}\}_{i=1}^{N}$ generated by an unknown behavior policy $\mu(\cdot|s)$, that can later generalize to any downstream task $z$ defined by a reward function $r_{\text{eval}}:\mathcal{S}\to\mathbb{R}$, i.e.
+The goal of the zero-shot RL problem is to train a task-agnostic policy given an offline dataset $\mathcal{D}=\{s_{i},a_{i},s_{i+1}\}_{i=1}^{N}$ generated by an unknown behavior policy $\mu(\cdot|s)$, that can later generalize to any downstream task $z$ defined by a reward function $r_{\text{eval}}:\mathcal{S}\to\mathbb{R}$, i.e.
 find $\max_{\pi}\ \mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{\text{eval}}\left(s_{t+1}\right)|s_{0},a_{0},\pi]$, relying only on a small set of reward-labeled samples with no further finetuning.
@@
 $Q_{r}^{\pi}(s,a)=\int_{s_{+}\in\mathcal{S}}M^{\pi}(s,a,ds_{+})\,r(s_{+}).$ (2)
-Given representation space $\mathbb{R}^{d}$, state distribution $\rho$, task vector $z\in\mathbb{R}^{d}$ and policy $\pi_{z}$ parameterized by $z$, FB decouples $M^{\pi_{z}}$ by a pair of forward representation $F:\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and backward representation $B:\mathcal{S}\to\mathbb{R}^{d}$, resulting optimal policy $\pi_{z}$ as:
+Given a representation space $\mathbb{R}^{d}$, a state distribution $\rho$, a task vector $z\in\mathbb{R}^{d}$ and a policy $\pi_{z}$ parameterized by $z$, FB expresses $M^{\pi_{z}}$ as the product of a forward representation $F:\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and a backward representation $B:\mathcal{S}\to\mathbb{R}^{d}$, resulting in an optimal policy $\pi_{z}$:
 $M^{\pi_{z}}(s_{0},a_{0},ds_{+})\approx F(s_{0},a_{0},z)^{\top}B(s_{+})\rho(ds_{+}),\quad\pi_{z}(s):=\arg\max_{a}F(s,a,z)^{\top}z,$ (3)
@@
 (4) during offline learning could introduce OOD extrapolation errors.
-To resolve this problem, they introduce additional CQL-style [33] regularizers on $M^{\pi_{z}}$ (MCFB) or $Q_{z}$ (VCFB) in $\mathcal{L}_{\text{FB}}$ for OOD regularization (i.e., push down the corresponding values on OOD actions and push up values evaluated on offline data samples).
+To resolve this problem, they introduce additional CQL-style [33] regularizers on $M^{\pi_{z}}$ (MCFB) or $Q_{z}$ (VCFB) in $\mathcal{L}_{\text{FB}}$ for OOD regularization (i.e., decrease values for OOD actions and increase those for data samples).
 3 Pitfalls of Exiting FB-Based Methods
@@
 FB offers an elegant theoretical foundation for zero-shot RL; however, our empirical findings reveal that FB-based methods often produce inconsistent and biased representations in practice.
-To illustrate this, we evaluate the $M^{\pi_{z}}$- and $Q_{z}$-value distributions derived from the learned $F$ and $B$ representations of the vanilla FB [58], the more recent MCFB [26], and our method on ExORL [70] benchmark with RND [5] datasets.
+To illustrate this, we evaluate the $M^{\pi_{z}}$- and $Q_{z}$-value distributions derived from the learned $F$ and $B$ representations of the vanilla FB [58], the more recent MCFB [26], and our method on ExORL benchmark [70] with RND [5] datasets.
 For each method, we use its $B$ representation to derive the task vector $z$ under default settings with a fixed number of transitions.
-The $M^{\pi_{z}}$ values are computed from episode-start state-action pairs $(s_{0},a_{0})\sim\rho_{0}$ and randomly sampled transitions $s_{+}\sim\mathcal{D}$, while $Q$-values are evaluated using policy-induced actions $a\sim\pi_{z}(s)$.
+The $M^{\pi_{z}}$ values are computed from episode-start state-action pairs $(s_{0},a_{0})\sim\rho_{0}$ and randomly sampled batch of transitions $s_{+}\sim\mathcal{D}$, while $Q$-values are evaluated using policy-induced actions $a\sim\pi_{z}(s)$.
 Results across additional environments and with method VCFB [26] are provided in Appendix D.3.
@@
 (b) Impact of $B$ Networks
-Figure 2: Investigations of the modeling architecture on the vanilla FB [58]. We utilize the peak observed in the training stage of each run as a proxy for the architecture’s capability ceiling.
+Figure 2: Investigation of the modeling architecture on the vanilla FB [58]. We utilize the peak observed in the training stage of each run as a proxy for the architecture’s capability ceiling.
 The challenges in estimating $M^{\pi_{z}}$- and $Q_{z}$-value can be primarily attributed to the limited supervisory signal in the vanilla FB learning objective (Eq.
@@
 $\begin{gathered}V_{\pi_{z}}(s,z):={\max_{\begin{subarray}{c}a\in A\\ \mathrm{s.t.}\,\mu(a|s)>0\end{subarray}}}F(s,a,z)^{\top}z,\end{gathered}$ (6)
-where $z$ is the task vector and $\mu(a|s)$ is the behavior policy in dataset $\mathcal{D}$.
-Compared with the traditional formulation of state-value function, this optimization problem aims to learn an optimal task-conditioned $V$-function solely from dataset samples without explicitly estimating $\mu(a|s)$.
+where $z$ is the task vector and $\mu(a|s)$ is the behavior policy in the dataset $\mathcal{D}$.
+Compared with the traditional formulation of the state-value function, this optimization problem aims to learn an optimal task-conditioned $V$-function solely from dataset samples without explicitly estimating $\mu(a|s)$.
 It can be solved through a class of value objectives according to different regularization formulations [67, 18, 30].
@@
 $\mathcal{L}_{V_{\pi_{z}}}=\mathbb{E}_{(s,a)\sim\mathcal{D},z\sim\mathcal{Z}}\left[L_{2}^{\tau}\left(F(s,a,z)^{\top}z-V_{\pi_{z}}(s,z)\right)\right],$ (7)
-where $L_{2}^{\tau}(u)=|\tau-\mathbb{I}(u>0)|u^{2}$, with $\tau>0.5$ serving as the expectile parameter.
+where $L_{2}^{\tau}(u)=|\tau-\mathbb{I}(u<0)|u^{2}$, with $\tau>0.5$ serving as the expectile parameter.
 By minimizing the above equations, we can obtain a well-regularized and nearly optimal state-value function $V_{\pi_{z}}$ for different tasks.
@@
 This formulation of $\mathcal{T}^{\pi}$ implicitly introduces regularization to the $Q$-function. Following above, we modified $\mathcal{L}_{F}$ in Eq. (5) as following $\mathcal{L}_{F\text{-reg}}$:
-$\mathcal{L}_{F\text{-reg}}=\mathbb{E}_{\begin{subarray}{c}(s,a,s^{\prime})\sim\mathcal{D}\end{subarray}}\left[\left(F(s,a,z)^{\top}z-B(s^{\prime})^{\top}\mathbb{E}_{D}[BB^{\top}]^{-1}z-\gamma V_{\pi_{z}}(s_{t+1},z)\right)^{2}\right].$
+$\mathcal{L}_{F\text{-reg}}=\mathbb{E}_{\begin{subarray}{c}(s,a,s^{\prime})\sim\mathcal{D}\end{subarray}}\left[\left(F(s,a,z)^{\top}z-B(s^{\prime})^{\top}\mathbb{E}_{D}[BB^{\top}]^{-1}z-\gamma V_{\pi_{z}}(s^{\prime},z)\right)^{2}\right].$
 By substituting the potentially unstable target $Q$-approximation $F(s_{t+1},\pi_{z}(s_{t+1}),z)^{\top}z$ with a more well-behaved state-value function $V_{\pi_{z}}$, we can greatly stabilize the learning of representation, as well as respond to optimality demands simultaneously.
@@
 Having introduced regularization for the value function, we now turn to policy learning, which prior work suggests can be equally or even more critical for final performance [43].
-A good policy should effectively leverage the high-quality behavior and further generalize near data distribution.
+A good policy should effectively leverage high-quality behavior and further generalize near data distribution.
 To ensure that the policy is well regularized by the dataset samples, we replace the policy objective in Eq.
@@
 To compute the solution of the weighted behavior cloning objective in Eq.
-(11) in practice, a key challenge arises: how to efficiently model and sample from potentially highly complex and diverse distributions distilled from a reweighted behavior policy?
+(11) in practice, a key challenge arises: how to efficiently model and sample from potentially highly complex and diverse distributions distilled from a re-weighted behavior policy?
 Direct weighted behavior cloning with a Gaussian policy often fails to capture the complex, multi-modal policy distributions [62, 20, 7], which is a demand for arbitrary task learning.
@@
 This combination of expressive generative modeling and value-based selection is crucial for achieving robust, high-performance zero-shot generalization.
-4.3 Expressivity Enhancement for Representations Modeling
+4.3 Expressivity Enhancement for Representation Modeling
 The effectiveness of any policy—particularly a highly expressive diffusion policy—in leveraging learned action weights hinges on accurate value estimation.
 Biased value estimates can degrade policy learning to simple imitation, failing to redistribute probability mass toward superior actions.
-To fully exploit the representational capacity of our diffusion policy, the underlying value representations must precisely capture complex task and dynamics relationships.
+To fully exploit the representational capacity of our diffusion policy, the underlying value representations must precisely capture complex task and dynamic relationships.
 We thus introduce enhanced network architectures for the forward ($F$) and backward ($B$) representations.
-Figure 3: The architectures in BREEZE
+Figure 3: Architecture of BREEZE
 As shown in Section 3, we empirically identify a pair of attention-based architectures that improve zero-shot performance over the original FB implementations.
@@
 The forward network encodes state-task and state-action pairs using two separate linear encoders, following Touati et al.
+Touati et al.
 [58].
@@
 These two feature sets correspond to semantically distinct concepts in the MDP.
-To accurately approximate the future occurrence measure $M(s,a,s^{\prime},z)$ in Eq.
+To accurately approximate the measure $M(s,a,s^{\prime},z)$ in Eq.
 3, the representation $F$ must integrate information from both behavioral and task contexts, capturing their interdependencies effectively.
-We therefore model the two encoded features as a length-2 sequence and process them through self-attention blocks.
+We therefore model the two encoded features as a length-2 embedding sequence and process them through self-attention blocks.
 This allows bidirectional feature refinement between task conditions and agent behaviors.
-The resulting representation is projected to a $d$-dimensional space via a linear layer.
+The resulting representation is projected onto a $d$-dimensional space via linear layers.
 Backward Network.
@@
 We equip $B$ with a stack of standard transformer-based networks with multi-head attention.
-The final output is projected to a $d$-dimensional space, consistent with the forward network.
+The final output is projected onto a $d$-dimensional space, consistent with the forward network.
 These architecturally enhanced representations enable BREEZE to model complex relationships more accurately, leading to improved value estimates and policy performance.
 Further implementation details are provided in Appendix B.4.
-Table 1: IQM results on ExORL[70] benchamrk. We report the best overall performance when all tasks perform well. Each value is averaged over 10 rollouts on 5 random seeds.
+Table 1: IQM results on ExORL benchmark. We report the best overall performance when all tasks perform well. Each value is averaged over 10 rollouts on 5 random seeds.
 Dataset Domain SF-LAP FB VCFB MCFB HILP BREEZE
@@
 Quadruped $295\pm 46$ $288\pm 48$ $286\pm 34$ $446\pm 78$
-Table 3: Normalized scores on Kitchen environment
+Figure 4: Normalized scores on Kitchen environment
 5 Experiments
@@
 Environmental setup.
-Our main experiments are conducted on the ExORL [70] benchmark, which provides a variety of datasets collected by several unsupervised RL algorithms [34].
-We select datasets collected by 4 algorithms: RND [5], APS [38], DIAYN [11], PROTO [69].
+Our main experiments are conducted on the ExORL benchmark [70], which provides a variety of datasets collected by several unsupervised RL algorithms [34].
+We select datasets collected by 4 algorithms: RND [5], APS [38], DIAYN [11], and PROTO [69].
 The experiments span 3 domains and 12 tasks (Walker: Stand, Walk, Run, Flip; Jaco: Reach Top/Bottom Left/Right; Quadruped: Stand, Walk, Run, Jump), bringing the total to 48 state-based complex tasks for performance calculation.
-In addition, we consider four challenging multi-stage tasks in the D4RL [14] Franka Kitchen [19] domain with two datasets (mixed and partial), which require long-horizon sequential manipulation on 4 subtasks.
-The overall setups on two benchmarks involve 2 locomotion domains and 2 manipulation (goal-reaching) domains.
+In addition, we consider four challenging multi-stage tasks in the D4RL [14] Franka Kitchen domain [19] with two datasets (mixed and partial), which require long-horizon sequential manipulation on 4 subtasks.
+The overall setup involves 2 locomotion domains and 2 manipulation (goal-reaching) domains.
 For all goal-conditioned domains, we use the backward representation to calculate the goal features as the task vector $z$.
@@
 We only consider offline algorithms that could perform a zero-shot policy generalization for our evaluation.
-Our baselines includes: 1) SF-LAP [3, 4]: successor feature-based method with basic features to be Laplacian Eigenfunction (LAP) [64], 2) vanilla FB [57], 3) VCFB and MCFB [26]: FB-based offline algorithms with CQL [33]-style regularizer on successor measure $M^{\pi}$ and $Q$-value separately for OOD issue avoidance, 4) HILP [44]: a state-based representation modeling algorithm keeping distance-preserving temporal structures in latent space having zero-shot capability.
+Our baselines include: 1) SF-LAP [3, 4]: successor feature-based method with basic features to be Laplacian Eigenfunction (LAP) [64], 2) vanilla FB [57], 3) VCFB and MCFB [26]: FB-based offline algorithms with CQL-style regularizer [33] on successor measure $M^{\pi}$ and $Q$-value separately for OOD issue avoidance, 4) HILP [44]: a state-based representation modeling algorithm keeping distance-preserving temporal structures in latent space having zero-shot capability.
 Evaluation.
@@
 [26].
-We record the Interquartile Mean (IQM), a robust metric that reduces the influence of outliers, in the training stage.
+We record the Interquartile Mean (IQM), a robust metric that reduces the influence of outliers.
 We evaluate the checkpoints every 10,000 updates with 10 rollouts for each experiment.
@@
 We analyze the findings of our experiments through the following key questions:
-  • Does BREEZE possess zero-shot capability to novel tasks?
+  • Does BREEZE possess zero-shot capability for novel tasks?
 Following previous studies [1, 26], we record the highest aggregated mean performance across random seeds in the offline pretrain stage.
-Table 3 reports this score across all tasks on each domain of ExORL, averaged on 5 random seeds, with standard deviations.
+Table 4.3 reports this score across all tasks on each domain of ExORL, averaged on 5 random seeds, with standard deviations.
 BREEZE achieves the best or near-best returns in most domains, demonstrating superior zero-shot generalization capability.
 Moreover, BREEZE can enhance vanilla FB’s generalization capability in long-horizon tasks.
-As shown in Table 3, we draw the box with the performance reported in the previous study [44], and record BREEZE’s top averaged returns across 4 random seeds.
+As shown in Figure 4, we draw the box with the performance reported in the previous study [44], and record BREEZE’s top averaged returns across 4 random seeds.
 While vanilla FB struggles in these long-horizon tasks, BREEZE achieves significantly higher performance, suggesting its ability to better leverage the dataset for complex sequential manipulation.
@@
   • How does BREEZE perform facing different quality, diversity of datasets?
-Table 3 shows results on 100,000-transition subsets of ExORL datasets.
+Table 4.3 shows results on 100,000-transition subsets of ExORL datasets.
 BREEZE maintains a pronounced advantage over FB-based baselines in this small-sample regime.
@@
 Figure 5: Learning Curves on ExORL RND. The solid lines represent the average return over 5 random seeds, while the shaded area denotes the standard deviation.
-Figure 5: Learning Curves on 100k-subsample transitions on ExORL APS Walker. The solid lines represent the average return over 3 random seeds, while the shaded area denotes the standard deviation.
+Figure 6: Learning Curves on 100k-subsample transitions on ExORL APS Walker domain.
+The solid lines represent the average return over 3 random seeds, while the shaded area denotes the standard deviation.
 Dataset w/o FB Enhancement w/o Diffusion BREEZE
@@
 BREEZE is ablated using: a gaussian policy variant (w/o Diffusion) and the baseline architecture variant (w/o FB Enhancement).
-We process transitions with data normalization for the diffusion policy, as in a commonly used preprocessing step in previous diffusion-based methods, and cancel this process for backing to the default policy setting to have a reasonable comparison on the components’ contributions.
+We process transitions with data normalization for the diffusion policy, as in a commonly used preprocessing step in previous diffusion-based methods.
+We cancel this process and return to the default policy setting to have a reasonable comparison of the components’ contributions.
 As shown in Table 3, each component somehow individually slightly boosts performance, but their synergy creates a leap in zero-shot capability.
@@
 Other approaches relax the linearity assumption by modeling SR as full distributions using generative models [25, 55, 63, 12].
-Among these, FB has emerged as a key framework, reducing the reward dependency by learning the successor measures $M$ instead of $Q$ or $V$, and improving tractability by factorizing $M$ into a forward and backward components.
-Extensions like MCFB/VCFB [26] build on offline RL adding value constraints to reduce the overestimation from OOD rollouts.
-FB-CPR [56] introduces regularizes the $Q$-estimation by imitating pre-collected datasets in an online setting.
+Among all, FB has emerged as a pivotal framework, factorizing successor measures into forward and backward components and avoiding the potential collapse issue in SF at the same time.
+Extensions like MCFB/VCFB [26] build on offline RL, enforcing explicit value constraints to mitigate the over-estimation issues that arise from updating on OOD rollouts.
+FB-CPR [56] introduces a regularization of the $Q$-estimation by imitating pre-collected datasets in an online setting.
 FB-AWARE [6] introduces an autoregressive mechanism to handle the over-linearization problem in vanilla FB.
@@
 Offline RL.
-Numerous studies address offline RL, in which learning is restricted to a fixed dataset without online interaction, leading to distributional shift.
-The earliest approaches use policy constraints, enforcing the learned policy to stay close to behavior policy [17, 32, 15, 35] or the distribution coverage to stay near the dataset [36, 8].
+Numerous studies address offline RL, in which learning is restricted to a fixed dataset without online interaction, leading to a distributional shift.
+The earliest approaches use policy constraints, enforcing the learned policy to stay close to the behavior policy [17, 32, 15, 35] or the distribution coverage to stay near the dataset [36, 8].
 Another established line of work employs value regularization, which directly penalizes the value estimates of OOD actions [33, 31, 42, 68, 66].
@@
 Diffusion models [48, 22, 37], known for their strong multi-modal modeling capability, have been well studied in imitation and trajectory modeling [24, 73, 9, 41, 52], and have recently been integrated into offline RL frameworks.
-Some approaches use the value function as guidance for sampling [39, 40].
+Some approaches use the value function as energy guidance for sampling [39, 40].
 Other works, such as SfBC [7], IDQL [20], and EDP [27], extract a diffusion policy via weighted regression under the in-sample learning framework.
@@
 In this work, we present BREEZE, a novel framework that mitigates the improper scaling issue in existing FB-based zero-shot RL methods.
-Specifically, BREEZE tackles two critical limitations: offline extrapolation errors and constrained policy expressivity.
+Specifically, BREEZE tackles two critical limitations: offline extrapolation errors and constrained expressivity.
 Our solution integrates behavior-regularized value estimation with a task-conditioned diffusion policy, enabling stable in-sample learning while capturing complex, multimodal action distributions.
@@
 Extensive experiments demonstrate that BREEZE consistently outperforms existing methods across various benchmarks, underscoring the importance of calibrated regularization and sufficient model capacity for zero-shot generalization.
-While BREEZE demonstrates improved robustness, its primary limitation is the increased computational cost from diffusion-based sampling—a common trade-off for high-performance generative policies.
+While BREEZE demonstrates improved robustness, its primary limitation is the increased computational cost from diffusion-based sampling, which is a common trade-off for utilizing high-performance generative policies.
 Further discussion is provided in Appendix E.
@@
 This work is supported by the Wuxi Research Institute of Applied Technologies, Tsinghua University, under Grant 20242001120, and funded by Horizon Robotics, AsiaInfo, and the Xiongan AI Institute.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Yes, the paper discusses its limitations in Appendix E, including the high computational demand.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: Yes, the paper provides the full set of assumptions and complete proofs for all theorems in the appendix A.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: Yes, the paper fully discloses all information needed to reproduce the main experimental results.
-Pseudocode, all experimental settings, hyperparameters, and baseline details are included in the appendix B-C.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Yes, the code will be made available after the open-source approval process is completed.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Yes, all hyperparameters and experimental setup details necessary to understand the results are provided in the appendix B.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: Yes, for all algorithms, several random seeds are used. All results are presented as mean ± standard deviation.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Yes, the Appendix C and E provide detailed information about the specific computational devices used for the experiments.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We make sure to preserve anonymity.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: Yes, the main paper discusses both potential positive societal impacts and limitations of the work.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification:
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Yes, all sources of data and code are properly credited in the appendix B, with licenses and terms of use clearly indicated.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification:
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer:[N/A]
-  69. Justification:
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification:
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
 Appendix A Theoretical Proofs
@@
 $z$ inference steps 10,000
-Temperature $\alpha$ 0.05 (0.1 kitchen-mixed)
 Regularization weight for orthogonality loss $w_{1}$ 10
 Beta schedule vp
+Exponential advantages clip $(-\infty,100]$
 Architecture $F$ preprocessor hidden dimension 1024
@@
 Table 5: Domain-specific hyperparameters for BREEZE with full dataset.
-Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$
-Walker-RND 9 0.1 64 0.99 0.001 5
-Walker-APS 9 0.1 64 0.99 0.001 5
-Walker-PROTO 9 0.1 64 0.99 0.001 5
-Walker-DIAYN 9 0.1 64 0.99 0.001 5
-Jaco-RND 1 0 64 0.99 0.001 10
-Jaco-APS 9 0.1 64 0.99 0.0001 5
-Jaco-PROTO 1 0 64 0.99 0.001 10
-Jaco-DIAYN 9 0.1 64 0.99 0.0001 5
-Quadruped-RND 1 0 64 0.99 0.001 10
-Quadruped-APS 1 0 64 0.99 0.001 10
-Quadruped-PROTO 1 0 64 0.99 0.001 10
-Quadruped-DIAYN 1 0 64 0.99 0.001 10
-Kitchen-mixed 1 0 16 0.7 0.001 5
-Kitchen-partial 2 0.2 8 0.7 0.001 5
+Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$ Temperature $\alpha$
+Walker-RND 9 0.1 64 0.99 0.001 5 0.05
+Walker-APS 9 0.1 64 0.99 0.001 5 0.05
+Walker-PROTO 9 0.1 64 0.99 0.001 5 0.05
+Walker-DIAYN 9 0.1 64 0.99 0.001 5 0.05
+Jaco-RND 1 0 64 0.99 0.001 10 0.05
+Jaco-APS 9 0.1 64 0.99 0.0001 5 0.05
+Jaco-PROTO 1 0 64 0.99 0.001 10 0.05
+Jaco-DIAYN 9 0.1 64 0.99 0.0001 5 0.05
+Quadruped-RND 1 0 64 0.99 0.001 10 0.05
+Quadruped-APS 1 0 64 0.99 0.001 10 0.05
+Quadruped-PROTO 1 0 64 0.99 0.001 10 0.05
+Quadruped-DIAYN 1 0 64 0.99 0.001 10 0.05
+Kitchen-mixed 1 0 16 0.7 0.001 5 0.1
+Kitchen-partial 2 0.2 4 0.7 0.001 5 0.08
 Table 6: Domain-specific hyperparameters for BREEZE with small sample dataset.
-Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$
-Walker-RND 2 0.2 32 0.99 0.0001 10
-Walker-APS 2 0.2 32 0.99 0.0001 10
-Walker-PROTO 2 0.2 32 0.99 0.001 10
-Walker-DIAYN 8 0.2 32 0.99 0.001 10
-Jaco-RND 1 0 32 0.99 0.0001 10
-Jaco-APS 1 0 32 0.99 0.0001 10
-Jaco-PROTO 1 0 32 0.99 0.0001 10
-Jaco-DIAYN 1 0 32 0.99 0.0001 10
-Quadruped-RND 2 0.2 32 0.99 0.001 10
-Quadruped-APS 1 0 32 0.99 0.001 10
-Quadruped-PROTO 2 0.2 32 0.99 0.001 10
-Quadruped-DIAYN 2 0.2 32 0.99 0.0001 10
+Domain-dataset $K_{\text{train}}$ $\rho_{a}$ $K_{\text{eval}}$ expectile $\tau$ $F\text{-reg}$ coef. $\omega_{q}$ Diffusion steps $T$ Temperature $\alpha$
+Walker-RND 2 0.2 32 0.99 0.0001 10 0.05
+Walker-APS 2 0.2 32 0.99 0.0001 10 0.05
+Walker-PROTO 2 0.2 32 0.99 0.001 10 0.05
+Walker-DIAYN 8 0.2 32 0.99 0.001 10 0.05
+Jaco-RND 1 0 32 0.99 0.0001 10 0.05
+Jaco-APS 1 0 32 0.99 0.0001 10 0.05
+Jaco-PROTO 1 0 32 0.99 0.0001 10 0.05
+Jaco-DIAYN 1 0 32 0.99 0.0001 10 0.05
+Quadruped-RND 2 0.2 32 0.99 0.001 10 0.05
+Quadruped-APS 1 0 32 0.99 0.001 10 0.05
+Quadruped-PROTO 2 0.2 32 0.99 0.001 10 0.05
+Quadruped-DIAYN 2 0.2 32 0.99 0.0001 10 0.05
 Appendix C Pseudo-Code
@@
 This section presents the zero-shot learning curves on the ExORL benchmark.
-Results on the full dataset and the 100k small-sample dataset are shown in Figure 9-11 and Figure 12-14, respectively.
+Results on the full dataset and the 100k small-sample dataset are shown in Figure 10-12 and Figure 13-15, respectively.
 The solid lines represent the mean IQM returns over 5 random seeds, and the shaded regions correspond to the standard deviation.
@@
 Evaluations were conducted every 10,000 steps during offline training.
-Following standard zero-shot RL criteria, we computed the aggregate score across all tasks as the overall performance measure, and selected the highest domain-level score across 5 random seeds along with corresponding per-task results.
+We computed the aggregate score across all tasks as the overall performance measure, and selected the highest domain-level score across 5 random seeds along with corresponding per-task results.
 Table 7: Full dataset experimental results on ExORL
@@
 D.3 Empirically Evaluated $M^{\pi_{z}}$ and $Q_{z}$ on Other Tasks
-We present empirical visualization results of $M^{\pi_{z}}$ and $Q_{z}$ distributions, shown in Figure 18-18.
+We present empirical visualization results of $M^{\pi_{z}}$ and $Q_{z}$ distributions, shown in Figure 19-19.
 We trained FB, VCFB, and MCFB on the RND [5] dataset.